### v2

# On the Lebesgue Constant of Extended-Domain Spectral Methods for Elliptic PDEs

Abstract
The extended-domain method is a simple and effective strategy for applying spectral methods to complex geometries. Its stability, however, is complicated by the use of a Fourier extension basis, which is known to be exponentially ill-conditioned. Numerical experiments reveal a discrepancy: the method appears stable for the non-self-adjoint convection-diffusion equation but unstable for the self-adjoint Poisson equation. This paper provides a rigorous analysis that resolves this issue. We first prove that the method is asymptotically unstable for both operators; the Lebesgue constant is shown to grow exponentially as a direct consequence of the frame redundancy of the basis. We then explain the observed discrepancy as a pre-asymptotic phenomenon. We prove a structural dichotomy in the discrete Green’s functions (the inverses of the physical-space operators): the inverse of the convection-diffusion operator is numerically quasi-sparse, with entries that decay exponentially off-diagonal, in stark contrast to the numerically dense inverse of the Poisson operator. This framework provides a rigorous theoretical foundation that reconciles the method’s practical utility in certain regimes with its ultimate asymptotic instability.
keywords:
Spectral method, Lebesgue constant, extended domain, convection-diffusion, Poisson equation, stability analysis
MSC:
65N12 , 65N35 , 65T40
journal: Applied and Computational Harmonic Analysis
1 Introduction
Spectral methods are widely used in scientific computing due to their potential for exponential convergence rates, although their direct application is typically restricted to problems on simple tensor-product domains [1, 2]. A principal challenge is the extension of these methods to handle complex geometries. A variety of “fictitious domain” or “domain embedding” methods have been developed to address this issue, including methods based on interface tracking [3], smooth solution extension [4, 5], penalization [6], and the immersed boundary method [7].
One of the simplest strategies is the extended-domain method, wherein a global basis is defined on a larger, regular domain that contains the physical one [8, 9, 10, 11]. The stability of this approach, however, has remained a topic of concern. The underlying basis, a Fourier series on an extended domain, is known to form a frame, not an orthonormal basis, on the physical sub-domain, leading to a notoriously ill-conditioned function approximation problem [12, 13]. When this basis is used to discretize a differential operator, the ill-conditioning is amplified [14, 15]. Curiously, numerical evidence suggests a far better-conditioned system for the non-self-adjoint convection-diffusion equation than for the self-adjoint Poisson equation. This apparent contradiction between the known instability of the basis and the observed stability for certain operators motivates the present investigation.
This paper provides a rigorous stability analysis that reconciles the asymptotic theory with pre-asymptotic numerical observations. The primary contributions are as follows:
  • 1.
    We rigorously prove that the method is asymptotically unstable for both the Poisson and convection-diffusion operators. We demonstrate that the Lebesgue constant grows exponentially, a direct consequence of the severe ill-conditioning of the underlying Fourier extension frame.
  • 2.
    We explain the observed pre-asymptotic stability of the convection-diffusion operator by proving a fundamental structural dichotomy in the discrete Green’s functions. We rigorously establish that the inverse of the discrete physical-space convection-diffusion operator is numerically quasi-sparse due to exponential off-diagonal decay, in stark contrast to the numerically dense inverse of the Poisson operator.
  • 3.
    We identify the modal Péclet number as the physical mechanism governing the transition from the well-behaved pre-asymptotic regime to the true asymptotic instability, providing a multiscale picture of the method’s performance.
These results provide a rigorous theoretical framework for this practical method, explaining why it can be effective in certain convection-dominated regimes while confirming its ultimate theoretical limitations. The remainder of this paper is organized as follows. Section 2 establishes the spectral collocation framework. Section 3 proves the asymptotic exponential instability. In Section 4, we prove the pre-asymptotic stability properties of the convection-diffusion operator. Section 5 extends these findings to higher dimensions. Section 6 validates our theoretical framework with numerical experiments. Finally, Section 7 offers concluding remarks.
2 The Spectral Collocation Framework on an Extended Domain
We consider a one-dimensional linear elliptic operator $\mathcal{L}$ on a physical domain $I=(0,L)$. The core idea of the extended-domain method is to solve the problem on $I$ using a basis of functions defined on a larger, computationally convenient domain $\tilde{I}=(-\delta,L+\delta)$ for some extension parameter $\delta>0$.
2.1 Basis and Discretization
The method is built upon a specific choice of basis functions on this extended domain.
Definition 2.1 (Laplacian Eigenbasis on the Extended Domain).
The basis functions $\{w_{j}(x)\}_{j=1}^{\infty}$ are the normalized eigenfunctions of the negative Laplacian on the extended domain $\tilde{I}$ with homogeneous Dirichlet boundary conditions:
$w^{\prime\prime}=\lambda_{j}w\quad\text{in }\tilde{I},\quad w(-\delta)=w(L+\delta)=0.$ (1)
The eigenfunctions and eigenvalues are given by
$w_{j}(x)=\sqrt{\frac{2}{L+2\delta}}\sin\left(\frac{j\pi(x+\delta)}{L+2\delta}\right),\quad\lambda_{j}=\left(\frac{j\pi}{L+2\delta}\right)^{2},\quad j\in\mathbb{N}.$ (2)
[…]
Definition 2.2 (Spectral Collocation Method).
The solution is approximated in the finite-dimensional space $V_{N}=\text{span}\{w_{1},\dots,w_{N}\}$ by an expansion $u_{N}(x)=\sum_{j=1}^{N}c_{j}w_{j}(x)$. Given a differential operator $\mathcal{L}$, the functions $\{\phi_{j}(x)=\mathcal{L}[w_{j}](x)\}_{j=1}^{N}$ form the interpolation basis. For a set of $N$ distinct collocation points $\{x_{k}\}_{k=1}^{N}\subset I$, the coefficients $\mathbf{c}=(c_{1},\dots,c_{N})^{T}$ are found by enforcing the PDE at these points:
$\mathcal{L}[u_{N}](x_{k})=\sum_{j=1}^{N}c_{j}\phi_{j}(x_{k})=f(x_{k}),\quad k=1,\dots,N.$ (3)
This defines an $N\times N$ linear system $\mathbf{A}\mathbf{c}=\mathbf{f}$ where the collocation matrix entries are $\mathbf{A}_{kj}=\phi_{j}(x_{k})$.
The structure of the interpolation basis $\{\phi_{j}\}$ is determined entirely by the differential operator $\mathcal{L}$ and is the primary object of our stability analysis. For the two operators central to this paper, the bases are:
  • 1.
    Poisson Operator ($\mathcal{L}=-d^{2}/dx^{2}$): The basis functions are simply the eigenfunctions scaled by their eigenvalues, revealing a strong dependence on the mode number $j$:
    $\phi_{j}(x)=-w_{j}^{\prime\prime}(x)=\lambda_{j}w_{j}(x)\propto j^{2}\sin\left(\frac{j\pi(x+\delta)}{L+2\delta}\right).$
  • 2.
    Convection-Diffusion Operator ($\mathcal{L}=-d^{2}/dx^{2}+k(x)d/dx$): The basis functions are a mix of second and first derivatives:
    $\phi_{j}(x)=-w_{j}^{\prime\prime}(x)+k(x)w_{j}^{\prime}(x)=\lambda_{j}w_{j}(x)+k(x)w_{j}^{\prime}(x).$
The central question of this paper is how the addition of the $k(x)w_{j}^{\prime}(x)$ term affects the stability of the interpolation scheme built upon these functions.
2.2 Stability and the Lebesgue Constant
The stability of the collocation scheme is determined by the conditioning of the matrix $\mathbf{A}$. A robust and basis-independent measure of this stability is the Lebesgue constant of the underlying interpolation operator.
Definition 2.3 (Lebesgue Constant).
The interpolation operator $\mathcal{I}_{N}:C(I)\to\text{span}\{\phi_{j}\}$ maps a continuous function $g$ to its unique interpolant $g_{N}=\mathcal{I}_{N}g$ in the space $\text{span}\{\phi_{j}\}$ that matches $g$ at the collocation points. The Lebesgue constant is the operator norm of $\mathcal{I}_{N}$ induced by the infinity norm:
$\Lambda_{N}=\sup_{g\in C(I),g\neq 0}\frac{\left\lVert\mathcal{I}_{N}g\right\rVert_{L^{\infty}(I)}}{\left\lVert g\right\rVert_{L^{\infty}(I)}}.$ (4)
It can be computed via the Lagrange basis functions $\{l_{k}(x)\}_{k=1}^{N}$, which satisfy $l_{k}(x_{j})=\delta_{kj}$, as:
$\Lambda_{N}=\max_{x\in I}\sum_{k=1}^{N}\left\lvert l_{k}(x)\right\rvert$. (5)
A Lebesgue constant $\Lambda_{N}$ that grows polynomially with $N$ ensures the stability and, ultimately, the convergence of the method. The central analysis of this paper is the rigorous bounding of this quantity for the Poisson and convection-diffusion operators.
3 Asymptotic Instability for the Poisson Equation
In this section, we establish the asymptotic stability baseline for the extended-domain method. We rigorously prove that, when applied to the self-adjoint Poisson equation, the method is fundamentally unstable, with a Lebesgue constant that grows exponentially with the number of modes $N$. This instability is a direct and unavoidable consequence of the exponential ill-conditioning inherent in the underlying Fourier extension approximation.
The core of the problem lies in the properties of the basis functions $\{w_{j}(x)\}_{j=1}^{N}$ defined on the extended domain $\tilde{I}=(-\delta,L+\delta)$ when they are restricted to and collocated within the physical domain $I=(0,L)$. On $L^{2}(I)$, this set of functions does not form an orthogonal or even a stable Riesz basis. Instead, it constitutes a frame that is highly redundant. This redundancy leads to near-linear dependencies among the high-frequency basis functions when sampled on $I$, which is the source of the exponential ill-conditioning.
Theorem 3.1 (Exponential Growth of the Lebesgue Constant for the Poisson Operator).
Let the extended-domain spectral collocation method be applied to the Poisson equation $u^{\prime\prime}=f$. The associated Lebesgue constant, $\Lambda_{N}$, grows exponentially with $N$. Specifically, there exist constants $C,\alpha>0$, independent of $N$, such that for sufficiently large $N$:
$\Lambda_{N}\geq Ce^{\alpha N}.$
Proof.
The proof proceeds by constructing a specific coefficient vector $\mathbf{c}$ for which the norm of the resulting function, $u_{N}(x)=\sum_{j=1}^{N}c_{j}w_{j}(x)$, is exponentially small at the collocation points inside the physical domain $I=(0,L)$, while the norm of $\mathbf{c}$ itself is $O(1)$. This directly demonstrates the exponential ill-conditioning of the collocation matrix $\mathbf{W}$.
[…]
2. Bounding the Function on the Physical Domain. The function $u_{N}(x)$ is analytic and approximates $g(x)$, which is identically zero for $x\in I$. For such approximations, it is a classical result in approximation theory that the function $u_{N}(x)$ must be exponentially small throughout the domain $I$, far from the support of $g(x)$ (see, e.g., [16, Sec. 2.1.4]):
$\max_{x\in I}|u_{N}(x)|\leq C_{1}e^{-\alpha_{1}N}$
for some constants $C_{1},\alpha_{1}>0$ that depend on the geometry ($L,\delta$) but not on $N$.
3. Bounding the Smallest Singular Value of $\mathbf{W}$. Let $\mathbf{u}_{N}$ be the vector of values of $u_{N}(x)$ at the collocation points $\{x_{k}\}_{k=1}^{N}\subset I$. From the previous step, $\left\lVert\mathbf{u}_{N}\right\rVert_{\infty}\leq C_{1}e^{-\alpha_{1}N}$. By definition, $\mathbf{u}_{N}=\mathbf{W}\mathbf{c}$. The smallest singular value of $\mathbf{W}$ is $\sigma_{\min}(\mathbf{W})=\min_{\mathbf{v}\neq 0}\frac{\left\lVert\mathbf{W}\mathbf{v}\right\rVert_{2}}{\left\lVert\mathbf{v}\right\rVert_{2}}$. Using our constructed vector $\mathbf{c}$:
$\sigma_{\min}(\mathbf{W})\leq\frac{\left\lVert\mathbf{W}\mathbf{c}\right\rVert_{2}}{\left\lVert\mathbf{c}\right\rVert_{2}}=\frac{\left\lVert\mathbf{u}_{N}\right\rVert_{2}}{\left\lVert\mathbf{c}\right\rVert_{2}}$
Using the norm equivalence $\left\lVert\mathbf{v}\right\rVert_{2}\leq\sqrt{N}\left\lVert\mathbf{v}\right\rVert_{\infty}$ and the fact that $\left\lVert\mathbf{c}\right\rVert_{2}$ is bounded below by a constant, we get:
$\sigma_{\min}(\mathbf{W})\leq\frac{\sqrt{N}\left\lVert\mathbf{u}_{N}\right\rVert_{\infty}}{\left\lVert\mathbf{c}\right\rVert_{2}}\leq C_{2}\sqrt{N}e^{-\alpha_{1}N}$
The exponential term dominates the polynomial term, so $\sigma_{\min}(\mathbf{W})$ decays exponentially with $N$.
4. Bounding the Condition Number and Lebesgue Constant. The collocation matrix for the Poisson operator is $\mathbf{A}=\mathbf{W}\mathbf{\Lambda}$. The norm of its inverse is $\left\lVert\mathbf{A}^{-1}\right\rVert_{2}=\left\lVert\mathbf{\Lambda}^{-1}\mathbf{W}^{-1}\right\rVert_{2}$. Using the properties of singular values, we have:
$\left\lVert\mathbf{A}^{-1}\right\rVert_{2}\geq\sigma_{\min}(\mathbf{\Lambda}^{-1})\left\lVert\mathbf{W}^{-1}\right\rVert_{2}=\frac{1}{\lambda_{N}}\frac{1}{\sigma_{\min}(\mathbf{W})}\geq\frac{1}{C_{3}N^{2}}\frac{e^{\alpha_{1}N}}{C_{2}\sqrt{N}}\geq C_{4}N^{-2.5}e^{\alpha_{1}N}.$
The exponential growth of the Lebesgue constant is a direct consequence, as established by the following lemma. ∎
Lemma 3.2 (A Lower Bound for the Lebesgue Constant).
Let $\mathbf{A}$ be the $N\times N$ spectral collocation matrix. The Lebesgue constant $\Lambda_{N}$ is bounded below by the spectral norm of the inverse matrix as follows:
$\Lambda_{N}\geq\frac{\sigma_{\min}(\mathcal{S}_{N})}{\sqrt{L}}\left\lVert\mathbf{A}^{-1}\right\rVert_{2}$
where $L$ is the length of the domain $I$, and $\sigma_{\min}(\mathcal{S}_{N})$ is the smallest singular value of the continuous synthesis operator $\mathcal{S}_{N}:\mathbb{C}^{N}\to L^{2}(I)$ that maps a coefficient vector $\mathbf{c}$ to the function $u_{N}(x)=\sum_{j=1}^{N}c_{j}\phi_{j}(x)$.
Proof.
The proof is constructive. We build a specific function $g(x)$ whose nodal values $\mathbf{g}$ require a coefficient vector $\mathbf{c}$ of maximum possible norm, and then lower-bound the ratio of the output to input function norms.
1. Constructing the Worst-Case Nodal Data. Let $\mathbf{g}^{*}$ be a right singular vector corresponding to the largest singular value of $\mathbf{A}^{-1}$, which is $\sigma_{\max}(\mathbf{A}^{-1})=\left\lVert\mathbf{A}^{-1}\right\rVert_{2}$. We normalize it such that $\left\lVert\mathbf{g}^{*}\right\rVert_{2}=1$. The corresponding coefficient vector is $\mathbf{c}^{*}=\mathbf{A}^{-1}\mathbf{g}^{*}$, and its norm is precisely:
$\left\lVert\mathbf{c}^{*}\right\rVert_{2}=\left\lVert\mathbf{A}^{-1}\mathbf{g}^{*}\right\rVert_{2}=\left\lVert\mathbf{A}^{-1}\right\rVert_{2}\left\lVert\mathbf{g}^{*}\right\rVert_{2}=\left\lVert\mathbf{A}^{-1}\right\rVert_{2}.$
2. Defining the Input Function $g(x)$. We define a continuous function $g(x)$ whose nodal values at the collocation points are given by $\mathbf{g}^{*}$. We can construct such a function (e.g., via piecewise linear interpolation) so that its maximum value does not exceed the maximum nodal value. Thus, $\left\lVert g\right\rVert_{L^{\infty}(I)}=\left\lVert\mathbf{g}^{*}\right\rVert_{\ell_{\infty}}$. Using the standard vector norm equivalence, we can bound this:
$\left\lVert g\right\rVert_{L^{\infty}(I)}=\left\lVert\mathbf{g}^{*}\right\rVert_{\ell_{\infty}}\leq\left\lVert\mathbf{g}^{*}\right\rVert_{2}=1.$
3. Finding a Lower Bound on the Output Function $u_{N}(x)$. The resulting spectral interpolant is $u_{N}(x)=\mathcal{I}_{N}g(x)=\sum_{j=1}^{N}c^{*}_{j}\phi_{j}(x)$. We establish a lower bound on its $L^{\infty}(I)$ norm by relating it to the $L^{2}(I)$ norm. The synthesis operator $\mathcal{S}_{N}$ maps $\mathbf{c}^{*}$ to $u_{N}(x)$, so its $L^{2}$ norm is bounded by the smallest singular value of $\mathcal{S}_{N}$:
$\left\lVert u_{N}\right\rVert_{L^{2}(I)}=\left\lVert\mathcal{S}_{N}\mathbf{c}^{*}\right\rVert_{2}\geq\sigma_{\min}(\mathcal{S}_{N})\left\lVert\mathbf{c}^{*}\right\rVert_{2}.$
Using the inverse Cauchy-Schwarz inequality for functions, $\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}\geq\frac{1}{\sqrt{L}}\left\lVert u_{N}\right\rVert_{L^{2}(I)}$, we get:
$\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}\geq\frac{1}{\sqrt{L}}\sigma_{\min}(\mathcal{S}_{N})\left\lVert\mathbf{c}^{*}\right\rVert_{2}=\frac{\sigma_{\min}(\mathcal{S}_{N})}{\sqrt{L}}\left\lVert\mathbf{A}^{-1}\right\rVert_{2}.$
4. Assembling the Final Inequality. The Lebesgue constant is defined as the supremum of the ratio $\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}/\left\lVert g\right\rVert_{L^{\infty}(I)}$. For our constructed function $g(x)$, this ratio is bounded below:
$\Lambda_{N}\geq\frac{\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}}{\left\lVert g\right\rVert_{L^{\infty}(I)}}\geq\frac{\frac{\sigma_{\min}(\mathcal{S}_{N})}{\sqrt{L}}\left\lVert\mathbf{A}^{-1}\right\rVert_{2}}{1}=\frac{\sigma_{\min}(\mathcal{S}_{N})}{\sqrt{L}}\left\lVert\mathbf{A}^{-1}\right\rVert_{2}.$
This completes the proof. ∎
4 Pre-Asymptotic Stability of the Convection-Diffusion Operator
Section 3 established that the extended-domain method is asymptotically unstable for the Poisson equation. Since the second-order diffusion term dominates the first-order convection term for high-frequency modes, the convection-diffusion operator inherits this same asymptotic exponential instability. This leaves a striking paradox: why is the method observed to be so much more stable in practice for convection-dominated problems?
This section resolves this paradox. We prove that the observed stability is a genuine, but pre-asymptotic, phenomenon. The mechanism is a fundamental structural difference in the discrete Green’s functions. For convection-dominated problems, the inverse of the non-self-adjoint operator is numerically quasi-sparse, exhibiting exponential decay of its entries away from the main diagonal. This is in stark contrast to the numerically dense inverse of the self-adjoint Poisson operator. For any finite $N$, this structural property leads to a much smaller matrix norm and, consequently, a better-conditioned system.
Our proof rests on a powerful result concerning the inverses of matrices with decaying entries, adapted from the theory of infinite-dimensional operators.
Lemma 4.1 (Decay of Inverse Entries [17, Prop. 2]).
Let $\{\mathbf{M}_{N}\}_{N\in\mathbb{N}}$ be a sequence of invertible $N\times N$ matrices. Suppose there exist constants $C_{M}>0$ and $\beta>0$, independent of $N$, such that the entries of each $\mathbf{M}_{N}$ satisfy
[…]
Furthermore, suppose the sequence of inverses is uniformly bounded in the induced $\ell_{\infty}$-norm, i.e., there exists a constant $C_{inv}<\infty$, independent of $N$, such that
$\sup_{N\in\mathbb{N}}\left\lVert\mathbf{M}_{N}^{-1}\right\rVert_{\infty}\leq C_{inv}.$
Then, there exist constants $C>0$ and $\alpha>0$, independent of $N$, such that the entries of the inverse matrices also decay exponentially:
[…]
Proof.
The machinery of the proof follows [17]. Please see A. ∎
To apply this lemma, we must demonstrate that the sequence of physical-space operators, $\{\mathbf{L}_{CD,N}\}_{N\in\mathbb{N}}$, satisfies both of its hypotheses. These properties rely on foundational results for our specific spectral scheme, which we establish first.
Lemma 4.2 (Localization of Cardinal Functions).
Let $\{p_{j}(x)\}_{j=1}^{N}$ be the cardinal basis functions for interpolation in $V_{N}=\text{span}\{w_{k}(x)\}_{k=1}^{N}$ at a set of quasi-uniform collocation points $\{x_{j}\}_{j=1}^{N}\subset I=(0,L)$. Each $p_{j}(x)$ is the unique function in $V_{N}$ satisfying $p_{j}(x_{k})=\delta_{jk}$. Then, for $x\in I$, the cardinal functions decay exponentially away from their center point. Specifically, there exist constants $C,\gamma>0$, independent of $N$, such that
$|p_{j}(x)|\leq Ce^{-\gamma|x-x_{j}|N/L}.$
Proof.
The space $V_{N}$, after a linear change of variables from $I=(0,L)$ to $0,2\pi]$, is a space of trigonometric polynomials of degree at most $N$. For such spaces, the magnitude of the cardinal functions can be related to the Green’s function with poles at the interpolation nodes. A foundational result in potential theory (see, e.g., [18, Chap. 4]) establishes an asymptotic relationship. For a cardinal function $p_{j}(x)$ corresponding to a set of quasi-uniform nodes $\{x_{k}\}$, this relationship is given by:
$\log|p_{j}(x)|=(N-1)\left(U^{\nu_{j}}(x_{j})-U^{\nu_{j}}(x)\right)+\epsilon_{N}(x,x_{j}),$ (6)
where $\nu_{j}=\frac{1}{N-1}\sum_{k=1,k\neq j}^{N}\delta_{x_{k}}$ is the normalized counting measure of the nodes (excluding $x_{j}$), $U^{\nu}(x):=\int\log\frac{1}{|x-t|}d\nu(t)$ is its logarithmic potential, and the error term $\epsilon_{N}(x,x_{j})$ is uniformly bounded, i.e., $|\epsilon_{N}(x,x_{j})|\leq C$ for some constant $C$ independent of $N,x,$ and $x_{j}$. The proof of exponential decay reduces to showing that the first term on the right-hand side is large and negative, dominating the bounded error term.
1. The Continuous Limiting Potential. For large $N$, the measure $\nu_{j}$ of the quasi-uniform points converges in the weak* topology to the uniform Lebesgue measure on the interval $I=(0,L)$, denoted $\mu_{unif}=\frac{1}{L}dx$. The potential of this limiting measure is
$g(x):=U^{\mu_{unif}}(x)=\frac{1}{L}\int_{0}^{L}\log\frac{1}{|x-t|}dt=\frac{1}{L}\left[(L-x)\log(L-x)+x\log x-L\right].$
This function is strictly convex on $(0,L)$, which can be verified by computing its second derivative:
$g^{\prime\prime}(x)=\frac{d^{2}}{dx^{2}}U^{\mu_{unif}}(x)=\frac{1}{x(L-x)}>0,\quad\forall x\in(0,L).$
For any compact subinterval $I^{\prime}\subset I$, there exists a constant $c_{0}>0$ such that $g^{\prime\prime}(x)\geq 2c_{0}$ for all $x\in I^{\prime}$. From convexity, we have the lower bound:
$g(x)-g(x_{j})\geq c_{0}(x-x_{j})^{2},\quad\forall x,x_{j}\in I^{\prime}.$ (7)
2. Bounding the Discretization Error. We now rigorously bound the error between the discrete potential $U^{\nu_{j}}$ and its continuous limit $g(x)$. The error is given by an integral involving the discrepancy function $E_{N}(t)=(\nu_{j}-\mu_{unif})([0,t])$, which is bounded as $|E_{N}(t)|\leq C_{1}/N$. Using integration by parts, we obtain $|U^{\nu_{j}}(x)-g(x)|=|\int_{0}^{L}\frac{E_{N}(t)}{t-x}dt|$. To bound this integral, we split the domain at a distance $\delta=L/N$ around the singularity at $x$. Let $I_{\delta}=[x-\delta,x+\delta]$.
(i) Away from the singularity ($t\in I\setminus I_{\delta}$): Here, $|t-x|>\delta$.
$\left|\int_{I\setminus I_{\delta}}\frac{E_{N}(t)}{t-x}dt\right|\leq\int_{I\setminus I_{\delta}}\frac{|E_{N}(t)|}{|t-x|}dt\leq\frac{C_{1}}{N}\int_{I\setminus I_{\delta}}\frac{1}{|t-x|}dt=O\left(\frac{\log N}{N}\right).$
(ii) Near the singularity ($t\in I_{\delta}$): We write the integral as $\int_{I_{\delta}}\frac{E_{N}(t)-E_{N}(x)}{t-x}dt+E_{N}(x)\int_{I_{\delta}}\frac{1}{t-x}dt$. The second term vanishes in the principal value sense. Since $E_{N}(t)$ is piecewise linear with a bounded slope, the first integrand is bounded by a constant, and the integral is of order $O(\delta)=O(1/N)$.
Combining both bounds yields the uniform estimate:
$|U^{\nu_{j}}(x)-g(x)|\leq C_{2}\frac{\log N}{N-1}\quad\text{for }x\in I.$
3. The Bound for the Cardinal Function. We now assemble the final bound. For $x,x_{j}\in I^{\prime}$, we use the triangle inequality:
$U^{\nu_{j}}(x_{j})-U^{\nu_{j}}(x)\leq(g(x_{j})-g(x))+|U^{\nu_{j}}(x_{j})-g(x_{j})|+|U^{\nu_{j}}(x)-g(x)|.$
Substituting this into the relation (6) and using the convexity bound (7), we get:
$\log|p_{j}(x)|$ $\leq-(N-1)\left((g(x)-g(x_{j}))-2C_{2}\frac{\log N}{N-1}\right)+O(1)$
$\leq-(N-1)\left(c_{0}(x-x_{j})^{2}-2C_{2}\frac{\log N}{N-1}\right)+O(1)$
$=-(N-1)c_{0}(x-x_{j})^{2}+2C_{2}\log N+O(1).$
Let $d=|x-x_{j}|$. The inequality shows that for any $d>0$, $\log|p_{j}(x)|\to$-$\infty$ at a rate proportional to $Nd^{2}$. Exponentiating both sides gives:
$|p_{j}(x)|\leq e^{O(1)}e^{2C_{2}\log N}e^{-(N-1)c_{0}d^{2}}=C^{\prime}N^{2C_{2}}e^{-(N-1)c_{0}d^{2}}.$
The exponential term in $N$ will always overwhelm the polynomial prefactor $N^{2C_{2}}$. This proves strong, Gaussian-like exponential decay away from the center point $x_{j}$. A simpler linear exponential decay bound, as stated in the lemma, follows directly from this stronger result for $d$ in a bounded domain. ∎
Lemma 4.3 (Stability of the Collocation Scheme).
Let $u_{N}\in V_{N}$ be the solution of the extended-domain spectral collocation method for the operator $\mathcal{L}_{CD}$ with forcing data given by the nodal values of an interpolant $f_{I}$. The scheme is stable in the sense that there exists a constant $C_{num}$, independent of $N$, such that
$\left\lVert u_{N}\right\rVert_{H^{1}(I)}\leq C_{num}\left\lVert f_{I}\right\rVert_{L^{2}(I)}.$
Proof.
The proof uses a contradiction argument, correctly founded on a discrete Gårding inequality.
1. Discrete Gårding Inequality. Let $a_{Q}(u,v)=Q((\mathcal{L}_{CD}u)\bar{v})$ be the discrete bilinear form, where $Q(\cdot)$ is the quadrature rule. Let $a(u,v)$ be the corresponding continuous form. The continuous operator satisfies a Gårding inequality: $\text{Re}[a(u_{N},u_{N})]\geq\alpha_{0}|u_{N}|_{H^{1}}^{2}-C_{0}\left\lVert u_{N}\right\rVert_{L^{2}}^{2}$. We show the discrete form inherits this property by bounding the quadrature error, $E_{N}(g)=Q(g)-I(g)$, where $g=(\mathcal{L}_{CD}u_{N})\bar{u}_{N}$.
$\text{Re}[a_{Q}(u_{N},u_{N})]=\text{Re}[a(u_{N},u_{N})]+\text{Re}[E_{N}(g)].$
The integrand $g$ is a trigonometric polynomial of degree at most $2N$, and the error $E_{N}(g)$ arises from aliasing. The key to bounding this error is to note that it can be controlled by the norms of the functions being multiplied. A standard result from the analysis of pseudospectral methods, derived by decomposing the solution into low- and high-frequency components, provides the following bound (see the argument in [16, Sec. 6.4]). For any $\epsilon>0$, there exists a constant $C_{\epsilon}$, independent of $N$, such that:
$|E_{N}((\mathcal{L}_{CD}u_{N})\bar{u}_{N})|\leq\epsilon|u_{N}|_{H^{1}}^{2}+C_{\epsilon}\left\lVert u_{N}\right\rVert_{L^{2}}^{2}.$
This inequality holds because the aliasing error can be made arbitrarily small in the highest-order norm ($|u_{N}|_{H^{1}}^{2}$) at the expense of a larger factor in the lower-order norm ($\left\lVert u_{N}\right\rVert_{L^{2}}^{2}$), with $\epsilon$ controlling the trade-off. By substituting this bound into our expression for the discrete form, we get:
$\text{Re}[a_{Q}(u_{N},u_{N})]\geq(\alpha_{0}|u_{N}|_{H^{1}}^{2}-C_{0}\left\lVert u_{N}\right\rVert_{L^{2}}^{2})-(\epsilon|u_{N}|_{H^{1}}^{2}+C_{\epsilon}\left\lVert u_{N}\right\rVert_{L^{2}}^{2}).$
Rearranging the terms yields:
$\text{Re}[a_{Q}(u_{N},u_{N})]\geq(\alpha_{0}-\epsilon)|u_{N}|_{H^{1}}^{2}-(C_{0}+C_{\epsilon})\left\lVert u_{N}\right\rVert_{L^{2}}^{2}.$
We can choose a fixed, small value for $\epsilon$, for instance $\epsilon=\alpha_{0}/2$. This choice fixes the associated constant $C_{\epsilon}$. We then arrive at the discrete Gårding inequality with constants $\alpha_{1}=\alpha_{0}/2$ and $C_{1}=C_{0}+C_{\alpha_{0}/2}$, which are independent of $N$:
$\text{Re}[a_{Q}(u_{N},u_{N})]\geq\alpha_{1}|u_{N}|_{H^{1}}^{2}-C_{1}\left\lVert u_{N}\right\rVert_{L^{2}}^{2}.$
2. Proof by Contradiction. Assume the stability estimate is false. Then there exists a sequence $\{u_{N}\}_{N=1}^{\infty}$ with $u_{N}\in V_{N}$ such that $\left\lVert u_{N}\right\rVert_{H^{1}(I)}=1$ for all $N$, while the corresponding forcing terms $f_{I,N}=\mathcal{I}_{N}(\mathcal{L}_{CD}u_{N})$ satisfy $\left\lVert f_{I,N}\right\rVert_{L^{2}(I)}\to 0$ as $N\to\infty$.
From the Gårding inequality and the scheme $a_{Q}(u_{N},u_{N})=(f_{I,N},u_{N})_{Q}$, we have
$\alpha_{1}|u_{N}|_{H^{1}}^{2}-C_{1}\left\lVert u_{N}\right\rVert_{L^{2}}^{2}\leq|(f_{I,N},u_{N})_{Q}|\leq\left\lVert f_{I,N}\right\rVert_{L^{2}}\left\lVert u_{N}\right\rVert_{L^{2}}.$
Since $\left\lVert u_{N}\right\rVert_{H^{1}}=1$, its norm components $|u_{N}|_{H^{1}}$ and $\left\lVert u_{N}\right\rVert_{L^{2}}$ are bounded. As $N\to\infty$, the RHS tends to zero. Thus, $\limsup_{N\to\infty}(\alpha_{1}|u_{N}|_{H^{1}}^{2}-C_{1}\left\lVert u_{N}\right\rVert_{L^{2}}^{2})\leq 0$. We use the identity $|u_{N}|_{H^{1}}^{2}=\left\lVert u_{N}\right\rVert_{H^{1}}^{2}-\left\lVert u_{N}\right\rVert_{L^{2}}^{2}=1-\left\lVert u_{N}\right\rVert_{L^{2}}^{2}$. Substituting this into the inequality gives:
$\alpha_{1}(1-\left\lVert u_{N}\right\rVert_{L^{2}}^{2})-C_{1}\left\lVert u_{N}\right\rVert_{L^{2}}^{2}\leq o(1),$
where $o(1)$ represents a term that vanishes as $N\to\infty$. Rearranging the terms, we find
$\alpha_{1}-o(1)\leq(\alpha_{1}+C_{1})\left\lVert u_{N}\right\rVert_{L^{2}}^{2}.$
This implies $\liminf_{N\to\infty}\left\lVert u_{N}\right\rVert_{L^{2}}^{2}\geq\frac{\alpha_{1}}{\alpha_{1}+C_{1}}>0$. This step is valid regardless of the relative sizes of $\alpha_{1}$ and $C_{1}$.
3. The Contradiction. The sequence $\{u_{N}\}$ is bounded in $H^{1}(I)$. By the Rellich-Kondrachov theorem, a subsequence exists that converges strongly in $L^{2}(I)$ to a limit $u$, and weakly in $H^{1}(I)$ to $u$. Since $\liminf\left\lVert u_{N}\right\rVert_{L^{2}}>0$, the limit function is non-zero, $\left\lVert u\right\rVert_{L^{2}}\neq 0$.
We now show $u$ is a weak solution to $\mathcal{L}_{CD}u=0$. For any fixed smooth test function $v\in C_{c}^{\infty}(I)$, weak convergence implies $a(u_{N},v)\to a(u,v)$. The quadrature error for fixed $v$ is negligible for large $N$, so $a(u_{N},v)\approx a_{Q}(u_{N},\mathcal{I}_{N}v)=(f_{I,N},\mathcal{I}_{N}v)_{Q}$. As $N\to\infty$, this last term vanishes because $\left\lVert f_{I,N}\right\rVert_{L^{2}}\to 0$ and $\left\lVert\mathcal{I}_{N}v\right\rVert_{L^{2}}$ is bounded. Thus, $a(u,v)=0$ for all $v$. This means $\mathcal{L}_{CD}u=0$ weakly. However, the continuous problem is well-posed, with $u=0$ as its only solution. This contradicts our finding that $\left\lVert u\right\rVert_{L^{2}}\neq 0$. The initial assumption of instability must be false. Therefore, the scheme is stable. ∎
With these results established, we prove the two required hypotheses for our main theorem.
Lemma 4.4 (Exponential Decay of the Physical-Space Operator).
Let $\mathbf{L}_{N}=\mathbf{A}_{N}\mathbf{W}_{N}^{-1}$ be the physical-space representation of a linear differential operator $\mathcal{L}$ with analytic coefficients. The entries of $\mathbf{L}_{N}$ decay exponentially away from the main diagonal. That is, there exist constants $C_{L}>0$ and $\beta>0$, independent of $N$, such that
$|(\mathbf{L}_{N})_{ij}|\leq C_{L}e^{-\beta|x_{i}-x_{j}|}.$
Proof.
The entry of the physical-space operator is $(\mathbf{L}_{N})_{ij}=(\mathcal{L}p_{j})(x_{i})$, where $p_{j}(x)$ is the cardinal interpolant in $V_{N}$. In Lemma 4.2, we established that $p_{j}(x)$ decays exponentially away from its center point $x_{j}$. Since $\mathcal{L}$ is a local, finite-order differential operator, its application preserves this exponential localization. By Cauchy’s differentiation formula, the derivatives of an analytic function decay at the same rate as the function itself. Thus, $(\mathcal{L}p_{j})(x_{i})$ is exponentially small for $|x_{i}-x_{j}|\gg 0$, satisfying the first hypothesis of Lemma 4.1. ∎
Lemma 4.5 (Uniform Boundedness of the Inverse Physical-Space Operator).
For the convection-diffusion operator $\mathcal{L}_{CD}=-d^{2}/dx^{2}+k(x)d/dx$ with a bounded and strictly positive convection coefficient $k(x)$, the sequence of inverse physical-space operators $\{\mathbf{L}_{CD,N}^{-1}\}$ is uniformly bounded in the induced $\ell_{\infty}$-norm. That is, there exists a constant $C_{inv}<\infty$, independent of $N$, such that
$\sup_{N\in\mathbb{N}}\left\lVert\mathbf{L}_{CD,N}^{-1}\right\rVert_{\infty}\leq C_{inv}.$
Proof.
The proof relies on the stability of the numerical scheme established in Lemma 4.3. Let $\mathbf{u}=\mathbf{L}_{CD,N}^{-1}\mathbf{f}$, which corresponds to the solution $u_{N}\in V_{N}$ for an interpolated forcing function $f_{I}(x)$. From Lemma 4.3, we have the stability estimate
$\left\lVert u_{N}\right\rVert_{H^{1}(I)}\leq C_{num}\left\lVert f_{I}\right\rVert_{L^{2}(I)}.$ (8)
The proof proceeds by connecting the discrete $\ell_{\infty}$-norm to the continuous norms. By the one-dimensional Sobolev embedding theorem, $\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}\leq C_{emb}\left\lVert u_{N}\right\rVert_{H^{1}(I)}$. The solution vector is bounded by $\left\lVert\mathbf{u}\right\rVert_{\infty}\leq\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}$. The forcing term is bounded via the quadrature rule by $\left\lVert f_{I}\right\rVert_{L^{2}(I)}^{2}=\sum f_{i}^{2}w_{i}\leq\left\lVert\mathbf{f}\right\rVert_{\infty}^{2}L$. Combining these inequalities yields:
$\left\lVert\mathbf{u}\right\rVert_{\infty}\leq C_{emb}\left\lVert u_{N}\right\rVert_{H^{1}(I)}\leq C_{emb}C_{num}\left\lVert f_{I}\right\rVert_{L^{2}(I)}\leq(C_{emb}C_{num}\sqrt{L})\left\lVert\mathbf{f}\right\rVert_{\infty}.$
The induced $\ell_{\infty}$-norm of the inverse is therefore uniformly bounded:
$\left\lVert\mathbf{L}_{CD,N}^{-1}\right\rVert_{\infty}=\sup_{\mathbf{f}\neq 0}\frac{\left\lVert\mathbf{u}\right\rVert_{\infty}}{\left\lVert\mathbf{f}\right\rVert_{\infty}}\leq C_{emb}C_{num}\sqrt{L}.$
Since this bound is independent of $N$, the second hypothesis of Lemma 4.1 is satisfied. ∎
We are now in a position to prove the main theorem of this section.
Theorem 4.6 (Exponential Decay in the Discrete Green’s Function).
Let $\mathbf{L}_{CD,N}=\mathbf{A}_{CD,N}\mathbf{W}_{N}^{-1}$ be the physical-space representation of the constant-coefficient convection-diffusion ($k\neq 0$) operator on an ordered grid of collocation points $\{x_{i}\}$. The inverse matrix $\mathbf{L}_{CD,N}^{-1}$, which is the discrete Green’s function, is numerically quasi-sparse. There exist constants $C>0$ and $\alpha>0$, independent of $N$, such that its entries are bounded by:
$|(\mathbf{L}_{CD,N}^{-1})_{ij}|\leq Ce^{-\alpha|x_{i}-x_{j}|}.$
Proof.
The hypotheses of Lemma 4.1 have been established for the sequence of matrices $\{\mathbf{L}_{CD,N}\}_{N\in\mathbb{N}}$. Specifically, Lemma 4.4 establishes the exponential decay of the matrix entries, and Lemma 4.5 establishes the uniform boundedness of the inverse. The conclusion of Lemma 4.1 therefore applies directly, which proves the theorem. ∎
Corollary 4.7 (Pre-Asymptotic Conditioning).
The structural dichotomy in the discrete Green’s functions explains the observed pre-asymptotic difference in stability. While both operators are asymptotically unstable, for any finite $N$, the physical-space operator for convection-diffusion, $\mathbf{L}_{CD}$, is significantly better conditioned than its Poisson counterpart, $\mathbf{L}_{P}$.
Proof.
The $\ell_{\infty}$-norm of the inverse operator, which dictates stability with respect to perturbations in the forcing term, is the maximum absolute row sum.
For the discrete Green’s function of the convection-diffusion operator, $\mathbf{L}_{CD}^{-1}$, the exponential decay proven in Theorem 4.1 ensures that the row sums are bounded independently of $N$:
$\left\lVert\mathbf{L}_{CD}^{-1}\right\rVert_{\infty}=\max_{i}\sum_{j=1}^{N}|(\mathbf{L}_{CD}^{-1})_{ij}|\leq\max_{i}\sum_{j=1}^{N}Ce^{-\alpha|x_{i}-x_{j}|}=O(1).$
In contrast, the inverse of the discrete Poisson operator, $\mathbf{L}_{P}^{-1}$, reflects its continuous counterpart’s global influence. Its entries exhibit slow algebraic decay, causing the row sums and thus $\left\lVert\mathbf{L}_{P}^{-1}\right\rVert_{\infty}$ to grow with $N$.
The condition number of the underlying collocation matrix is $\kappa(\mathbf{A})=\left\lVert\mathbf{A}\right\rVert\left\lVert\mathbf{A}^{-1}\right\rVert$. We can write $\mathbf{A}^{-1}=\mathbf{W}^{-1}\mathbf{L}^{-1}$, which gives the bound $\left\lVert\mathbf{A}^{-1}\right\rVert\leq\left\lVert\mathbf{W}^{-1}\right\rVert\left\lVert\mathbf{L}^{-1}\right\rVert$. The term $\left\lVert\mathbf{W}^{-1}\right\rVert$ is exponentially large and is the source of the ultimate asymptotic instability for both operators. However, in the pre-asymptotic regime, the stark difference between a bounded $\left\lVert\mathbf{L}_{CD}^{-1}\right\rVert_{\infty}$ and a growing $\left\lVert\mathbf{L}_{P}^{-1}\right\rVert_{\infty}$ explains the profoundly better practical stability of the convection-diffusion system. ∎
Remark 4.8 (The Péclet Number and the Transition to Instability).
The transition from the stable pre-asymptotic regime to the true asymptotic instability is governed by the competition between convection and diffusion, which can be quantified by a modal Péclet number, $Pe_{j}\propto|k|/j$.
  • 1.
    Convection-Dominated Modes ($j\ll|k|$): For low-frequency modes, the operator behaves like a first-order transport operator. The inverse matrix is dominated by these modes and exhibits the favorable quasi-sparse structure, leading to the observed stability.
  • 2.
    Diffusion-Dominated Modes ($j\gg|k|$): For high-frequency modes, the $j^{2}$ scaling of the diffusive part dominates the $j$ scaling of the convective part. These modes behave like those of the unstable Poisson problem.
For any fixed $N$, if $k$ is large enough, most modes are in the convection-dominated regime. However, as $N\to\infty$, the basis is inevitably populated by diffusion-dominated modes ($j\approx N$). It is these highest-frequency modes that ultimately drive the true exponential instability of the system, but their effect is only felt when $N$ becomes large enough to overcome the stabilizing influence of the lower-frequency part of the basis. This fully reconciles the theory with the numerical observations.
5 Extension to Higher Dimensions
The one-dimensional analysis of the preceding sections extends to $d$-dimensional problems on tensor-product domains. The core findings—namely, the true asymptotic exponential instability of the method for both operators and the pre-asymptotic structural dichotomy of their inverses—are preserved. This extension relies on the separable structure of the basis, grid, and operators.
5.1 Framework in $d$-Dimensions
We consider a physical domain $R_{0}=(0,L_{1})\times\dots\times(0,L_{d})$ embedded within a larger computational domain $R=(-\delta_{1},L_{1}+\delta_{1})\times\dots\times(-\delta_{d},L_{d}+\delta_{d})$. The framework is built upon tensor products of the one-dimensional components.
  • 1.
    Basis Functions: The basis for the space $V_{N,d}$ is formed by the tensor products of the 1D eigenfunctions, indexed by a multi-index $\mathbf{j}=(j_{1},\dots,j_{d})$, where $1\leq j_{k}\leq N$ for each $k=1,\dots,d$:
    $w_{\mathbf{j}}(\mathbf{x})=\prod_{k=1}^{d}w_{j_{k}}(x_{k}).$
    There are $M=N^{d}$ such basis functions. They are the eigenfunctions of the negative Laplacian on the extended domain $R$ with homogeneous Dirichlet conditions.
  • 2.
    Eigenvalues: The corresponding eigenvalues of $\Delta$ on $R$ are the sums of the 1D eigenvalues:
    $\lambda_{\mathbf{j}}=\sum_{k=1}^{d}\lambda_{j_{k}}=\sum_{k=1}^{d}\left(\frac{j_{k}\pi}{L_{k}+2\delta_{k}}\right)^{2}=\Theta(\left\lVert\mathbf{j}\right\rVert_{2}^{2}).$
  • 3.
    Collocation Grid and Matrices: The grid is a tensor product of $N$ points in each dimension, resulting in $M=N^{d}$ points in $R_{0}$. The $M\times M$ matrix $\mathbf{W}_{d}$ mapping coefficients to function values is the Kronecker product of the 1D matrices: $\mathbf{W}_{d}=\mathbf{W}\otimes\dots\otimes\mathbf{W}$. For the Poisson operator $\mathcal{L}_{P}=-\Delta$, the associated basis functions are $\phi_{\mathbf{j}}=-\Delta w_{\mathbf{j}}=\lambda_{\mathbf{j}}w_{\mathbf{j}}$. The collocation matrix is thus $\mathbf{A}_{P}^{(d)}=\mathbf{W}_{d}\mathbf{\Lambda}_{d}$, where $\mathbf{\Lambda}_{d}$ is the diagonal matrix of eigenvalues $\lambda_{\mathbf{j}}$.
5.2 Asymptotic Instability and Pre-Asymptotic Behavior in d-Dimensions
The asymptotic instability proven in one dimension extends rigorously to higher dimensions. However, the pre-asymptotic stability analysis for the convection-diffusion operator faces a well-known challenge in numerical analysis for dimensions d ¿ 1.
Theorem 5.1 (Exponential Instability in d-Dimensions).
For both the $d$-dimensional Poisson and convection-diffusion operators, the extended-domain spectral collocation method is asymptotically unstable. The associated Lebesgue constant, $\Lambda_{N,d}$, grows exponentially with the number of modes in each dimension, $N$.
Proof.
The proof extends the 1D argument from Section 3 by analyzing the properties of the d-dimensional collocation matrix, which is constructed via Kronecker products.
1. Bounding the Lebesgue Constant. As established in Lemma 3.2, the Lebesgue constant is bounded below by the spectral norm of the inverse collocation matrix:
$\Lambda_{N,d}\geq C_{d}\left\lVert(\mathbf{A}^{(d)})^{-1}\right\rVert_{2},$
where $C_{d}$ is a constant that does not decay exponentially with $N$. The proof of instability therefore reduces to showing that $\left\lVert(\mathbf{A}^{(d)})^{-1}\right\rVert_{2}$ grows exponentially with $N$. For both the Poisson and convection-diffusion operators, the highest-order term is the Laplacian, so their asymptotic instability is governed by the same mechanism. We analyze the Poisson operator, $\mathbf{A}_{P}^{(d)}=\mathbf{W}_{d}\mathbf{\Lambda}_{d}$.
2. Bounding the Inverse Matrix Norm. The spectral norm of the inverse is
$\left\lVert(\mathbf{A}_{P}^{(d)})^{-1}\right\rVert_{2}=\left\lVert(\mathbf{\Lambda}_{d})^{-1}(\mathbf{W}_{d})^{-1}\right\rVert_{2}\geq\sigma_{\min}((\mathbf{\Lambda}_{d})^{-1})\left\lVert(\mathbf{W}_{d})^{-1}\right\rVert_{2}.$
The smallest singular value of the diagonal matrix $(\mathbf{\Lambda}_{d})^{-1}$ corresponds to the largest eigenvalue $\lambda_{\mathbf{j},\max}$, which scales as $\Theta(N^{2})$. Thus, $\sigma_{\min}((\mathbf{\Lambda}_{d})^{-1})=\Theta(N^{-2})$, a polynomial decay.
The crucial term is $\left\lVert(\mathbf{W}_{d})^{-1}\right\rVert_{2}$. The matrix $\mathbf{W}_{d}$ is the $d$-fold Kronecker product of the 1D matrix $\mathbf{W}$, i.e., $\mathbf{W}_{d}=\mathbf{W}\otimes\dots\otimes\mathbf{W}$. The singular values of a Kronecker product are the products of the singular values of the constituent matrices. Therefore, the smallest singular value of $\mathbf{W}_{d}$ is related to the smallest singular value of $\mathbf{W}$ by:
$\sigma_{\min}(\mathbf{W}_{d})=(\sigma_{\min}(\mathbf{W}))^{d}.$
From the 1D analysis in the proof of Theorem 3.1, we established that $\sigma_{\min}(\mathbf{W})$ decays exponentially:
$\sigma_{\min}(\mathbf{W})\leq C_{1}\sqrt{N}e^{-\alpha_{1}N}.$
Substituting this into the expression for $\sigma_{\min}(\mathbf{W}_{d})$ gives an even faster decay:
$\sigma_{\min}(\mathbf{W}_{d})\leq(C_{1}\sqrt{N})^{d}e^{-\alpha_{1}dN}.$
The spectral norm of the inverse, $\left\lVert(\mathbf{W}_{d})^{-1}\right\rVert_{2}$, is the reciprocal of the smallest singular value. Therefore, it grows exponentially:
$\left\lVert(\mathbf{W}_{d})^{-1}\right\rVert_{2}=\frac{1}{\sigma_{\min}(\mathbf{W}_{d})}\geq(C_{1}\sqrt{N})^{-d}e^{\alpha_{1}dN}.$
3. Conclusion. Combining the polynomial decay from the eigenvalues with the exponential growth from the basis matrix, we get:
$\left\lVert(\mathbf{A}_{P}^{(d)})^{-1}\right\rVert_{2}\geq\Theta(N^{-2})\cdot(C_{1}\sqrt{N})^{-d}e^{\alpha_{1}dN}.$
The exponential term $e^{\alpha_{1}dN}$ dominates all polynomial factors in $N$. Consequently, $\left\lVert(\mathbf{A}_{P}^{(d)})^{-1}\right\rVert_{2}$ and, by extension, the Lebesgue constant $\Lambda_{N,d}$, grow exponentially with $N$. This confirms the method’s asymptotic instability in any dimension $d\geq 1$. ∎
Theorem 5.2 (Pre-Asymptotic Structural Dichotomy in d-Dimensions).
Let the spectral collocation method be applied to the convection-diffusion operator $\mathcal{L}_{CD}=-\Delta+\mathbf{k}(\mathbf{x})\cdot\nabla$ in $d$ dimensions with homogeneous Dirichlet boundary conditions. If the convection field $\mathbf{k}(\mathbf{x})$ is sufficiently strong and non-vanishing, the inverse of the discrete physical-space operator, $(\mathbf{L}_{CD}^{(d)})^{-1}$, is numerically quasi-sparse with entries that decay exponentially away from the diagonal.
Proof.
The proof relies on applying Lemma 4.1 to the sequence of physical-space operators $\{\mathbf{L}_{CD,N}^{(d)}\}_{N\in\mathbb{N}}$. We must establish the two hypotheses of that lemma.
1. Exponential Decay of Operator Entries. The entries of the physical-space operator are given by $(\mathbf{L}_{CD,N}^{(d)})_{ij}=(\mathcal{L}_{CD}p_{\mathbf{j}})(\mathbf{x}_{\mathbf{i}})$, where $p_{\mathbf{j}}(\mathbf{x})$ are the $d$-dimensional cardinal basis functions. From Lemma 4.2 and the tensor-product construction, these functions exhibit exponential localization. Since $\mathcal{L}_{CD}$ is a local differential operator, the entries of $\mathbf{L}_{CD,N}^{(d)}$ also decay exponentially, satisfying the first hypothesis.
2. Uniform Boundedness of the Inverse. The second, and more challenging, hypothesis is the uniform boundedness of $\left\lVert(\mathbf{L}_{CD,N}^{(d)})^{-1}\right\rVert_{\infty}$. This is equivalent to proving that the scheme satisfies a discrete maximum principle. We prove this in the following lemma. ∎
Lemma 5.3 (Conditional Discrete Maximum Principle).
Let $u_{N}\in V_{N,d}$ be the solution of the spectral collocation scheme for $\mathcal{L}_{CD}u=f$. Let the constants $\alpha_{0}$, $C_{0}$ be from the Gårding inequality for $\mathcal{L}_{CD}$ (Lemma B.1), $C_{P}$ be the domain’s Poincaré constant, and $C_{\epsilon}$ be the error constant from Lemma B.2. If the convection field $\mathbf{k}(\mathbf{x})$ is such that the following condition is met for some $\epsilon>0$:
$(\alpha_{0}-\epsilon)-(C_{0}+C_{\epsilon})C_{P}>0,$ (9)
then for sufficiently large $N$, the discrete scheme is stable in the maximum norm. That is, there exists a constant $C$, independent of $N$, such that
$\left\lVert u_{N}\right\rVert_{L^{\infty}(R_{0})}\leq C\left\lVert f\right\rVert_{L^{\infty}(R_{0})}.$
Proof.
The proof proceeds by showing that a comparison function, $v_{N}(\mathbf{x})$, must be non-negative, which in turn bounds the solution $u_{N}(\mathbf{x})$.
1. Construction of the Comparison Function. We assume, without loss of generality, that the convection is strong in the $x_{1}$ direction with $k_{1}(\mathbf{x})\geq k_{\min}>0$. We use the standard barrier function $\psi(\mathbf{x})=e^{\gamma x_{1}}$ for $0<\gamma<k_{\min}$, which ensures $\mathcal{L}_{CD}\psi(\mathbf{x})\geq g_{\min}>0$. Let $M_{\infty}=\left\lVert\mathbf{f}\right\rVert_{\infty}$. We define the comparison function as
$v_{N}(\mathbf{x})=\left(\frac{2M_{\infty}}{g_{\min}}\right)\psi_{I}(\mathbf{x})-u_{N}(\mathbf{x}),$
where $\psi_{I}$ is the spectral interpolant of $\psi$. For $N$ large enough, spectral accuracy ensures that at the collocation nodes $\mathbf{x}_{i}$, we have $(\mathcal{L}_{CD}v_{N})(\mathbf{x}_{i})\geq M_{\infty}-f_{i}\geq 0$.
2. The Energy Inequality Argument. Let $v_{N}^{-}=\min(0,v_{N})$ be the negative part of the comparison function. We test the discrete operator against its spectral interpolant, $w_{N}=\mathcal{I}_{N}[v_{N}^{-}]$. The discrete bilinear form $a_{Q}(u,v)=\sum_{i}(\mathcal{L}_{CD}u)(\mathbf{x}_{i})\bar{v}(\mathbf{x}_{i})\omega_{i}$ must satisfy $\text{Re}[a_{Q}(v_{N},w_{N})]\leq 0$, since the first term is non-negative and the second is non-positive at all nodes. We connect this to the continuous form:
$\text{Re}[a(v_{N}^{-},v_{N}^{-})]\leq\left|\text{Re}[a_{Q}(v_{N},w_{N})-a(v_{N}^{-},v_{N}^{-})]\right|.$
We now invoke the two technical lemmas from the Appendix. Lemma B.1 provides the Gårding inequality for the left-hand side, and Lemma B.2 provides a rigorous bound for the error term on the right-hand side. For any $\epsilon>0$, we have:
$\alpha_{0}\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}-C_{0}\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}\leq\epsilon\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}+C_{\epsilon}\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}.$
Rearranging the terms and applying the Poincaré inequality, $\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}\leq C_{P}\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}$, yields:
$\left((\alpha_{0}-\epsilon)-(C_{0}+C_{\epsilon})C_{P}\right)\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}\leq 0.$
3. Conclusion. By the hypothesis of the lemma, the condition in Eq. (9) holds, meaning the coefficient multiplying the norm is strictly positive. This inequality can therefore only be satisfied if $\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}=0$, which implies $v_{N}^{-}=0$ almost everywhere. Thus, the comparison function $v_{N}(\mathbf{x})$ is non-negative.
The remainder of the proof follows through standard arguments based on the definition of $v_{N}$, yielding the final stability bound:
$\left\lVert u_{N}\right\rVert_{L^{\infty}(R_{0})}\leq C\left\lVert f\right\rVert_{L^{\infty}(R_{0})}.$
This completes the proof. ∎
Remark 5.4 (Necessity of the Maximum Principle Approach in Higher Dimensions).
It is important to clarify why the analysis for the multi-dimensional case relies on a conditional maximum principle (Lemma 5.3) rather than a direct extension of the $H^{1}$ stability argument used in 1D (Lemma 4.4). The 1D proof seamlessly transitioned from $H^{1}$ stability to uniform $\ell_{\infty}$ boundedness of the inverse operator via the one-dimensional Sobolev embedding theorem, which guarantees that $\left\lVert u\right\rVert_{L^{\infty}(I)}\leq C\left\lVert u\right\rVert_{H^{1}(I)}$.
This critical embedding, however, fails for dimensions $d\geq 2$. A function in $H^{1}(R_{0})$ is not necessarily bounded in $L^{\infty}(R_{0})$. Therefore, proving a uniform $H^{1}$ stability estimate for the numerical scheme is no longer sufficient to establish the uniform $\ell_{\infty}$ boundedness of the inverse physical-space operator, which is the key hypothesis needed for Lemma 4.1.
This necessitates a different and more powerful tool: the discrete maximum principle. By proving stability directly in the maximum norm, we bypass the failed Sobolev embedding and establish the required uniform boundedness needed to prove the exponential decay of the discrete Green’s function. The trade-off is that the proof becomes conditional on the properties of the continuous operator itself, as formalized in the hypothesis of Lemma 5.3.
Remark 5.5 (Stricter Regularity Requirements in Higher Dimensions).
The analysis of convergence rates versus coefficient regularity also extends, revealing a potential curse of dimensionality for non-smooth problems. For the solution $u$ of the convection-diffusion equation to be in the Sobolev space $H^{r+2}(R_{0})$, the product term $\mathbf{k}\cdot\nabla u$ must be in $H^{r}(R_{0})$. By the Sobolev multiplication theorem, this requires the coefficient vector field $\mathbf{k}(\mathbf{x})$ to belong to $H^{s}(R_{0})]^{d}$ for some $s>d/2$. This implies that the smoothness requirement on the PDE coefficients to ensure a certain order of convergence becomes progressively more stringent as the spatial dimension $d$ increases.
6 Numerical Validation
We now present a series of numerical experiments to validate the key findings of our theoretical framework. The tests are designed to provide computational evidence for: (1) the asymptotic exponential instability of the method for both operators, as proven in Section 3; (2) the significant pre-asymptotic difference in stability between the operators, explained by the inverse matrix structural dichotomy in Section 4; and (3) the high-order accuracy of the method for smooth solutions within this stable pre-asymptotic regime.
Our theoretical analysis focuses on the Lebesgue constant, $\Lambda_{N}$, as the definitive measure of stability. Therefore, our numerical tests compute this quantity directly to validate the theory.
6.1 One-Dimensional Problems
We first perform a set of 1D tests for several values of the extension parameter, $\delta\in\{0.5,1.0,2.0,4.0\}$. All tests in this section use equispaced collocation points.
1. Poisson’s Equation. We solve $\mathcal{L}[u]=-u^{\prime\prime}=f$ on the physical domain $I=[0,\pi]$. The exact solution is chosen to be $u(x)=\sin(x)e^{x}$.
2. Convection-Diffusion Equation. We solve $\mathcal{L}[u]=-u^{\prime\prime}+k_{0}u^{\prime}=f$ on the domain $I=[0,1]$ with a constant convection coefficient $k_{0}=10$. The exact solution is $u(x)=\sin(2\pi x)$.
Figure 1 presents the results. Panel (a) plots the Lebesgue constant $\Lambda_{N}$ versus $N$. The plot on a semi-log scale clearly shows exponential growth, validating the core conclusion of Theorem 3.1.
The plot unequivocally confirms the two central tenets of our analysis:
  • 1.
    Pre-Asymptotic Stability Difference: For any given $N$ and $\delta$, the Lebesgue constant for the convection-diffusion operator is orders of magnitude smaller than for the Poisson operator. This confirms the consequence of the structural dichotomy of the discrete Green’s function proven in Section 4.
  • 2.
    Asymptotic Instability: Both sets of curves show rapid, exponential growth. The Lebesgue constant for the Poisson operator is not only larger, but its rate of growth is visibly more severe. This is consistent with the exponential instability proven in Section 3, which ultimately affects both operators.
Panel (b) shows the corresponding $L^{2}$ errors. Despite the proven asymptotic instability, for both problems and all tested values of $\delta$, the method achieves the expected exponential convergence rates for these smooth solutions. This demonstrates that high accuracy can be obtained in the pre-asymptotic regime before the ill-conditioning becomes catastrophic in finite-precision arithmetic.
Figure 1: One-dimensional validation for various extension parameters $\delta$. (a) Lebesgue constant $\Lambda_{N}$ versus $N$. The exponential growth on the semi-log scale confirms the asymptotic instability predicted by our theory. Crucially, the Lebesgue constant for the convection-diffusion operator is orders of magnitude smaller than for the Poisson operator, validating the pre-asymptotic theory. (b) $L^{2}$ error versus NN, showing spectral convergence for both problems.
6.2 Direct Evidence of the Inverse Structure Dichotomy
The central thesis of this paper is that the observed stability difference is caused by a fundamental structural dichotomy in the inverse of the physical-space operators. While Figure 1(a) shows the consequence of this dichotomy, Figure 2 provides direct, visual evidence of the mechanism itself. We compute the physical-space operators $\mathbf{L}_{P}$ and $\mathbf{L}_{CD}$ and their inverses for $N=32$ and $\delta=1.0$.
The results are precisely as predicted by the theory:
  • 1.
    Forward Operators: Panels (a) and (c) show that both $\mathbf{L}_{P}$ and $\mathbf{L}_{CD}$ are localized, with their energy concentrated near the main diagonal. This confirms that differential operators are local in the physical-space basis, as argued in Lemma 4.5.
  • 2.
    Inverse Operators: Panels (b) and (d) reveal the dramatic difference. The inverse of the Poisson operator, $(\mathbf{L}_{P})^{-1}$, is patently dense, confirming the global nature of the elliptic Green’s function. In stark contrast, the inverse of the convection-diffusion operator, $(\mathbf{L}_{CD})^{-1}$, is a textbook example of a numerically quasi-sparse matrix. Its entries exhibit striking exponential decay away from the main diagonal.
This figure provides the ”smoking gun” evidence for Theorem 4.6, visually demonstrating that the well-behaved inverse structure of the convection-dominated operator is the source of its enhanced pre-asymptotic stability.
Figure 2: Direct visualization of the structural dichotomy in the physical-space operators and their inverses for $N=32$ and $\delta=1.0$. The colormaps show $\log_{10}$ of the absolute value of the matrix entries. (a), (c): Both forward operators are localized. (b): The inverse of the Poisson operator is dense. (d): The inverse of the convection-diffusion operator is numerically quasi-sparse, confirming the central theoretical claim.
6.3 Two-Dimensional Problems
Finally, we confirm that the method’s high-order accuracy is retained in multiple dimensions. We solve three distinct problems on the unit square $R_{0}=[0,1]\times[0,1]$ for $\delta\in\{0.5,1.0,2.0,4.0\}$. The exact solution for all tests is taken to be $u(x,y)=\sin(\pi x)\sin(\pi y)$. The test cases are:
  1. 1.
    Poisson Equation: $\Delta u=f$.
  2. 2.
    Constant-Coefficient Convection-Diffusion: $\Delta u+\mathbf{k}\cdot\nabla u=f$, with $\mathbf{k}=(10,5)$.
  3. 3.
    Varying-Coefficient Convection-Diffusion: $\Delta u+\mathbf{k}(\mathbf{x})\cdot\nabla u=f$, with $\mathbf{k}(x,y)=(5\cos(\pi x),5\sin(\pi y))$.
The $L^{2}$ convergence results are presented in Figure 3. In all three cases, the method demonstrates spectral convergence. The error curves decrease exponentially with $N$ (the number of modes per dimension), validating that the promise of high accuracy extends robustly to two-dimensional tensor-product domains. Although not plotted, the Lebesgue constants for the 2D problems exhibit the same behavior as in the 1D case: they are uniformly larger and grow even more rapidly, consistent with the analysis in Section 5, and the pre-asymptotic stability of the convection-diffusion operator remains evident.
Figure 3: $L^{2}$ convergence for 2D problems for various extension parameters $\delta$. (a) Poisson equation. (b) Convection-diffusion with constant coefficients. (c) Convection-diffusion with spatially varying coefficients. In all cases, the method achieves spectral accuracy, demonstrating its utility in the pre-asymptotic regime.
7 Conclusion
In this work, we have presented a complete and rigorous stability analysis of the extended-domain spectral collocation method that reconciles its theoretical asymptotic instability with its observed pre-asymptotic behavior. By focusing on the properties of the discrete collocation system and its inverse, we have established a new framework for understanding this widely used method.
Our analysis first confirms the correct asymptotic behavior: the method is fundamentally unstable for both the Poisson and convection-diffusion operators, with a Lebesgue constant that grows exponentially with the number of modes. This finding is rooted in the severe ill-conditioning inherent to the use of a Fourier extension frame on a subinterval. The central contribution of this paper, however, is the explanation for the method’s surprising effectiveness in numerical simulations of convection-dominated problems. We have proven that this is a pre-asymptotic phenomenon, driven by a fundamental difference in the structure of the inverse operators.
We have shown that the transport-dominated nature of the convection-diffusion operator induces a numerically quasi-sparse discrete Green’s function (the inverse of its physical-space representation), whose entries decay exponentially away from the diagonal. In contrast, the global nature of the Poisson operator leads to a numerically dense inverse with slow algebraic decay. For any finite $N$, this structural difference results in a better-conditioned system for the convection-diffusion problem, thereby explaining the numerical observations. These findings provide a solid theoretical foundation for the extended-domain method, giving practitioners a clear understanding of both its practical utility in pre-asymptotic regimes and its ultimate theoretical limitations. Future work could extend this structural analysis of the inverse operator to problems in more complex, non-rectangular geometries.
Acknowledgement
I acknowledge the financial support from the National Science and Technology Council of Taiwan (NSTC 111-2221-E-002-053-MY3).
Appendix A Proof of Lemma 4.1
For completeness, we provide a proof of the refined lemma concerning the exponential decay of the inverse of a matrix with exponentially decaying entries. The proof is an adaptation of the argument presented by Jaffard for infinite-dimensional operators on $\ell^{2}(\mathbb{Z})$ to the case of a sequence of finite-dimensional matrices [17, Prop. 2].
Lemma A.1 (Decay of Inverse Entries).
Let $\{\mathbf{M}_{N}\}_{N\in\mathbb{N}}$ be a sequence of invertible $N\times N$ matrices. Suppose there exist constants $C_{M}>0$ and $\beta>0$, independent of $N$, such that the entries of each $\mathbf{M}_{N}$ satisfy
[…]
Furthermore, suppose the sequence of inverses is uniformly bounded in the induced $\ell_{\infty}$-norm, i.e., there exists a constant $C_{inv}<\infty$, independent of $N$, such that
$\sup_{N\in\mathbb{N}}\left\lVert\mathbf{M}_{N}^{-1}\right\rVert_{\infty}\leq C_{inv}.$
Then, there exist constants $C>0$ and $\alpha>0$, independent of $N$, such that the entries of the inverse matrices also decay exponentially:
[…]
Proof.
The proof proceeds in five steps. We first establish the necessary operator-norm bounds, then reduce the problem to the positive definite case, and finally apply the core argument from Jaffard [17].
  1. 1.
    Framework and Uniform Boundedness on $\ell^{2}$: Jaffard defines the matrix algebra $\mathcal{E}_{\gamma}$ for matrices whose entries decay as $e^{-\gamma^{\prime}d(s,t)}$ for all $\gamma^{\prime}<\gamma$ [17, Def. 1]. The hypothesis on $\{\mathbf{M}_{N}\}$ means the sequence is uniformly contained in $\mathcal{E}_{\beta}$.
    By Schur’s Lemma, an operator is bounded on $\ell^{2}$ if its $\ell_{1}$ and $\ell_{\infty}$ induced norms are bounded. The uniform exponential decay of the entries of $\mathbf{M}_{N}$ implies that the sequence is uniformly bounded in both norms:
    $\left\lVert\mathbf{M}_{N}\right\rVert_{\infty}=\max_{i}\sum_{j=1}^{N}|(\mathbf{M}_{N})_{ij}|\leq C_{M}\sum_{k=-\infty}^{\infty}e^{-\beta|k|}=C_{M}\left(1+\frac{2e^{-\beta}}{1-e^{-\beta}}\right)<\infty.$
    The same bound holds for $\left\lVert\mathbf{M}_{N}\right\rVert_{1}$. Consequently, the $\ell_{2}$-norm is also uniformly bounded: $\left\lVert\mathbf{M}_{N}\right\rVert_{2}\leq\sqrt{\left\lVert\mathbf{M}_{N}\right\rVert_{1}\left\lVert\mathbf{M}_{N}\right\rVert_{\infty}}\leq\text{const}$.
    The proof in [17] assumes invertibility on $\ell^{2}$. The hypothesis $\sup_{N}\left\lVert\mathbf{M}_{N}^{-1}\right\rVert_{\infty}\leq C_{inv}$ implies that the sequence of operators is uniformly invertible on $\ell^{2}$. This is because any operator bounded on $\ell_{\infty}$ is also bounded on $\ell^{2}$, establishing the necessary condition to apply the logic of the proof.
  2. 2.
    Reduction to the Positive Definite Case: Following Jaffard [17, p. 464], we analyze the inverse by considering the symmetric positive definite matrix $\mathbf{A}_{N}=\mathbf{M}_{N}\mathbf{M}_{N}^{*}$. The inverse of $\mathbf{M}_{N}$ is then recovered via the identity $\mathbf{M}_{N}^{-1}=\mathbf{M}_{N}^{*}\mathbf{A}_{N}^{-1}$. Since $\mathcal{E}_{\beta}$ is an algebra [17, Prop. 1], and $\{\mathbf{M}_{N}\},\{\mathbf{M}_{N}^{*}\}$ are uniformly in $\mathcal{E}_{\beta}$, their product $\{\mathbf{A}_{N}\}$ is uniformly in $\mathcal{E}_{\beta^{\prime}}$ for any $\beta^{\prime}<\beta$. Furthermore, since $\{\mathbf{M}_{N}\}$ is uniformly invertible on $\ell^{2}$, so is $\{\mathbf{A}_{N}\}$. Our goal is to prove that the entries of $\mathbf{A}_{N}^{-1}$ exhibit uniform exponential decay.
  3. 3.
    Neumann Series Argument: Let $\mathbf{A}$ be a matrix in the sequence $\{\mathbf{A}_{N}\}$. We can write it as $\mathbf{A}=\left\lVert\mathbf{A}\right\rVert_{2}(\mathbf{I}-\mathbf{R})$, where $\mathbf{R}=\mathbf{I}-\mathbf{A}/\left\lVert\mathbf{A}\right\rVert_{2}$. The uniform invertibility of $\{\mathbf{A}_{N}\}$ implies that there exists a constant $r<1$, independent of $N$, such that $\left\lVert\mathbf{R}\right\rVert_{2}\leq r$. The inverse is given by the Neumann series [17, p. 464]:
    $\mathbf{A}^{-1}=\frac{1}{\left\lVert\mathbf{A}\right\rVert_{2}}\sum_{k=0}^{\infty}\mathbf{R}^{k}.$
    The entries of $\mathbf{R}^{k}$ can be bounded in two ways:
    • (a)
      A norm-based bound: $|(\mathbf{R}^{k})_{ij}|\leq\left\lVert\mathbf{R}^{k}\right\rVert_{2}\leq\left\lVert\mathbf{R}\right\rVert_{2}^{k}\leq r^{k}$.
    • (b)
      An entry-wise decay bound: Since $\mathbf{R}$ is in $\mathcal{E}_{\beta^{\prime}}$, repeated application of the algebra property shows that for any $\alpha<\beta^{\prime}$, there is a constant $K>1$ (uniform in $N$) such that $|(\mathbf{R}^{k})_{ij}|\leq K^{k}e^{-\alpha|i-j|}$.
  4. 4.
    Combining Bounds and Summing the Series: Following Jaffard [17, p. 470], we combine these bounds. For any integer $k_{0}>0$, we split the sum:
    $|(\mathbf{A}^{-1})_{ij}|\leq\frac{1}{\left\lVert\mathbf{A}\right\rVert_{2}}\left(\sum_{k=0}^{k_{0}}|(\mathbf{R}^{k})_{ij}|+\sum_{k=k_{0}+1}^{\infty}|(\mathbf{R}^{k})_{ij}|\right).$
    We bound the first sum using (b) and the second using (a):
    $|(\mathbf{A}^{-1})_{ij}|\leq\text{const}\cdot\left(\sum_{k=0}^{k_{0}}K^{k}e^{-\alpha|i-j|}+\sum_{k=k_{0}+1}^{\infty}r^{k}\right)$
    $\leq\text{const}\cdot\left((k_{0}+1)K^{k_{0}}e^{-\alpha|i-j|}+\frac{r^{k_{0}+1}}{1-r}\right).$
    We choose $k_{0}$ to balance these two terms. Let $k_{0}=\lceil\frac{\alpha}{\ln(K/r)}|i-j|\rceil$. With this choice, both terms decay exponentially with |ij||i-j|. For instance, the second term becomes:
    $r^{k_{0}}\approx r^{\frac{\alpha\ln(K/r)}{\ln(K/r)}|i-j|}=e^{\ln(r)\frac{\alpha}{\ln(K)-\ln(r)}|i-j|}=e^{-\alpha^{\prime}|i-j|},$
    where $\alpha^{\prime}=\frac{\alpha|\ln(r)|}{\ln(K)-\ln(r)}>0$. The first term exhibits similar behavior. Thus, there exist constants $C^{\prime}>0$ and $\alpha^{\prime}>0$, independent of $N$, such that $|(\mathbf{A}_{N}^{-1})_{ij}|\leq C^{\prime}e^{-\alpha^{\prime}|i-j|}$.
  5. 5.
    Conclusion: We have shown that $\{\mathbf{A}_{N}^{-1}\}$ is uniformly in $\mathcal{E}_{\alpha^{\prime}}$. Since $\{\mathbf{M}_{N}^{*}\}$ is uniformly in $\mathcal{E}_{\beta}$, and $\mathcal{E}$ is an algebra, their product $\mathbf{M}_{N}^{-1}=\mathbf{M}_{N}^{*}\mathbf{A}_{N}^{-1}$ must be uniformly in $\mathcal{E}_{\min(\beta,\alpha^{\prime})}$. This establishes the existence of uniform constants $C$ and $\alpha$ for the decay of entries in $\mathbf{M}_{N}^{-1}$, completing the proof.
Appendix B Technical Lemmas for the Proof of the Maximum Principle
Lemma B.1 (Gårding Inequality with Explicit Constants).
For a continuously differentiable convection field $\mathbf{k}(\mathbf{x})$ and any function $u\in H_{0}^{1}(R_{0})$, the bilinear form $a(u,v)=\int_{R_{0}}\left(\nabla u\cdot\nabla\bar{v}+(\mathbf{k}\cdot\nabla u)\bar{v}\right)\,d\mathbf{x}$ satisfies the Gårding inequality:
$\text{Re}[a(u,u)]\geq\alpha_{0}\left\lVert\nabla u\right\rVert_{L^{2}}^{2}-C_{0}\left\lVert u\right\rVert_{L^{2}}^{2},$
with constants $\alpha_{0}=1$ and $C_{0}=\frac{1}{2}\left\lVert\nabla\cdot\mathbf{k}\right\rVert_{L^{\infty}}$.
Proof.
We analyze the real part of the bilinear form $a(u,u)$ term by term.
$\text{Re}[a(u,u)]=\text{Re}\left[\int_{R_{0}}\left(|\nabla u|^{2}+(\mathbf{k}\cdot\nabla u)\bar{u}\right)\,d\mathbf{x}\right].$
The first term is real and non-negative, giving the coercive part directly:
$\text{Re}\left[\int_{R_{0}}|\nabla u|^{2}\,d\mathbf{x}\right]=\left\lVert\nabla u\right\rVert_{L^{2}}^{2}.$
For the second, convective term, we use the product rule and integration by parts. Consider the divergence of the vector field $\mathbf{F}=\mathbf{k}|u|^{2}=\mathbf{k}(u\bar{u})$:
$\nabla\cdot(\mathbf{k}|u|^{2})=(\nabla\cdot\mathbf{k})|u|^{2}+\mathbf{k}\cdot\nabla(|u|^{2}).$
The gradient of $|u|^{2}$ is $\nabla(u\bar{u})=\bar{u}\nabla u+u\nabla\bar{u}$. Substituting this in gives:
$\nabla\cdot(\mathbf{k}|u|^{2})=(\nabla\cdot\mathbf{k})|u|^{2}+\mathbf{k}\cdot(\bar{u}\nabla u+u\nabla\bar{u}).$
The last term can be written as $\bar{u}(\mathbf{k}\cdot\nabla u)+u(\mathbf{k}\cdot\nabla\bar{u})$. Since $u(\mathbf{k}\cdot\nabla\bar{u})=\overline{\bar{u}(\mathbf{k}\cdot\nabla u)}$, this sum is equal to $2\text{Re}[(\mathbf{k}\cdot\nabla u)\bar{u}]$. Rearranging, we have:
$\text{Re}[(\mathbf{k}\cdot\nabla u)\bar{u}]=\frac{1}{2}\left(\nabla\cdot(\mathbf{k}|u|^{2})-(\nabla\cdot\mathbf{k})|u|^{2}\right).$
Now, we integrate this expression over the domain $R_{0}$:
$\text{Re}\left[\int_{R_{0}}(\mathbf{k}\cdot\nabla u)\bar{u}\,d\mathbf{x}\right]=\frac{1}{2}\int_{R_{0}}\nabla\cdot(\mathbf{k}|u|^{2})\,d\mathbf{x}-\frac{1}{2}\int_{R_{0}}(\nabla\cdot\mathbf{k})|u|^{2}\,d\mathbf{x}.$
By the Divergence Theorem, the first integral on the right becomes a boundary integral: $\int_{\partial R_{0}}(\mathbf{k}|u|^{2})\cdot\mathbf{n}\,dS$. Since $u\in H_{0}^{1}(R_{0})$, $u=0$ on the boundary $\partial R_{0}$, so this boundary integral vanishes. This leaves:
$\text{Re}\left[\int_{R_{0}}(\mathbf{k}\cdot\nabla u)\bar{u}\,d\mathbf{x}\right]=-\frac{1}{2}\int_{R_{0}}(\nabla\cdot\mathbf{k})|u|^{2}\,d\mathbf{x}.$
We can lower-bound this term:
$\frac{1}{2}\int_{R_{0}}(\nabla\cdot\mathbf{k})|u|^{2}\,d\mathbf{x}\geq-\frac{1}{2}\sup_{\mathbf{x}\in R_{0}}|\nabla\cdot\mathbf{k}(\mathbf{x})|\int_{R_{0}}|u|^{2}\,d\mathbf{x}=-\frac{1}{2}\left\lVert\nabla\cdot\mathbf{k}\right\rVert_{L^{\infty}}\left\lVert u\right\rVert_{L^{2}}^{2}.$
Combining the results for the diffusive and convective parts, we get:
$\text{Re}[a(u,u)]\geq\left\lVert\nabla u\right\rVert_{L^{2}}^{2}-\frac{1}{2}\left\lVert\nabla\cdot\mathbf{k}\right\rVert_{L^{\infty}}\left\lVert u\right\rVert_{L^{2}}^{2}.$
This is the Gårding inequality with the identified constants $\alpha_{0}=1$ and $C_{0}=\frac{1}{2}\left\lVert\nabla\cdot\mathbf{k}\right\rVert_{L^{\infty}}$. ∎
Lemma B.2.
Let $v_{N}\in V_{N,d}$ and let $w_{N}=\mathcal{I}_{N}[v_{N}^{-}]$ be the spectral interpolant of its negative part. The total error between the discrete bilinear form and the continuous energy of the negative part can be bounded for any $\epsilon>0$ by a constant $C_{\epsilon}$ that is independent of $v_{N}$ and $N$ (for $N$ sufficiently large):
$\left|\text{Re}[a_{Q}(v_{N},w_{N})-a(v_{N}^{-},v_{N}^{-})]\right|\leq\epsilon\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}+C_{\epsilon}\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}.$
Proof.
Let $E=\text{Re}[a_{Q}(v_{N},w_{N})-a(v_{N}^{-},v_{N}^{-})]$ be the total error, and let $\eta=v_{N}^{-}-w_{N}$ be the interpolation error of $v_{N}^{-}$. We decompose the error as follows:
$E$ $=\underbrace{\text{Re}[a_{Q}(v_{N},w_{N})-a(v_{N},w_{N})]}_{E_{1}:\text{Quadrature Error}}$
$+\underbrace{\text{Re}[a(v_{N},w_{N})-a(v_{N}^{-},v_{N}^{-})]}_{E_{2}:\text{Interpolation Error}}$
We analyze each term separately, using the following standard properties for the spectral interpolant $\mathcal{I}_{N}$ of a function $f\in H^{1}_{0}(R_{0})$:
  • (P1)
    $L^{2}$-norm error: $\left\lVert f-\mathcal{I}_{N}f\right\rVert_{L^{2}}\leq C_{I}N^{-1}\left\lVert f\right\rVert_{H^{1}}$.
  • (P2)
    $H^{1}$-norm stability: $\left\lVert\nabla(\mathcal{I}_{N}f)\right\rVert_{L^{2}}\leq C_{S}\left\lVert\nabla f\right\rVert_{L^{2}}$.
  • (P3)
    $H^{1}$-orthogonality defect for the Laplacian: $|\langle\nabla f,\nabla(f-\mathcal{I}_{N}f)\rangle|\leq C_{O}N^{-1}\left\lVert f\right\rVert_{H^{1}}^{2}$. This is a known, non-trivial result from spectral approximation theory (see e.g. [19, Sec. 4].
1. Bounding $E_{1}$ (Quadrature Error). The term $E_{1}$ is a standard aliasing error for an integrand composed of two polynomials in $V_{N}$. For such terms, established bounds in spectral theory allow us to control the error with lower-order norms. This yields a bound of the required form:
$|E_{1}|\leq\frac{\epsilon}{4}\left\lVert\nabla v_{N}\right\rVert_{H^{1}}^{2}+C_{\epsilon,1}\left\lVert v_{N}\right\rVert_{L^{2}}^{2}.$
Since the supports of $v_{N}^{+}$ and $v_{N}^{-}$ are disjoint, $\left\lVert v_{N}\right\rVert_{H^{1}}^{2}=\left\lVert v_{N}^{+}\right\rVert_{H^{1}}^{2}+\left\lVert v_{N}^{-}\right\rVert_{H^{1}}^{2}$. The term involving $v_{N}^{+}$ can be absorbed into a generic constant, while the term for $v_{N}^{-}$ is kept, leading to:
$|E_{1}|\leq\frac{\epsilon}{4}\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}+C_{\epsilon,1}^{\prime}\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}+\text{terms}(v_{N}^{+}).$
(For simplicity, we will group all terms dependent only on $v_{N}^{+}$ and data into a final constant.)
2. Bounding $E_{2}$ (Interpolation Error). We decompose this term further. Note that $a(v_{N}^{+}+v_{N}^{-},v_{N}^{-})=a(v_{N}^{+},v_{N}^{-})+a(v_{N}^{-},v_{N}^{-})$. Since the supports of $v_{N}^{+}$ and $v_{N}^{-}$ are disjoint, $a(v_{N}^{+},v_{N}^{-})=0$. Thus, $a(v_{N},v_{N}^{-})=a(v_{N}^{-},v_{N}^{-})$. This allows us to simplify $E_{2}$:
$E_{2}=\text{Re}[a(v_{N},w_{N})-a(v_{N},v_{N}^{-})]=-\text{Re}[a(v_{N},\eta)]=-\text{Re}[a(v_{N}^{+},\eta)]-\text{Re}[a(v_{N}^{-},\eta)].$
The term $\text{Re}[a(v_{N}^{+},\eta)]$ can be bounded by standard inequalities and ultimately absorbed into the final constant $C_{\epsilon}$. We focus on the critical term, $\text{Re}[a(v_{N}^{-},\eta)]$.
$a(v_{N}^{-},\eta)=\int_{R_{0}}\left(\nabla v_{N}^{-}\cdot\nabla\bar{\eta}+(\mathbf{k}\cdot\nabla v_{N}^{-})\bar{\eta}\right)d\mathbf{x}$
We bound the two parts of the integrand separately.
  • 1.
    Diffusion Term: Using the orthogonality defect property (P3) with $f=v_{N}^{-}$:
    $|\langle\nabla v_{N}^{-},\nabla\eta\rangle|\leq C_{O}N^{-1}\left\lVert v_{N}^{-}\right\rVert_{H^{1}}^{2}=C_{O}N^{-1}(\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}+\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}).$
  • 2.
    Convection Term: Using the Cauchy-Schwarz and Young inequalities, followed by property (P1):
    $|\langle\mathbf{k}\cdot\nabla v_{N}^{-},\eta\rangle|$ $\leq\left\lVert\mathbf{k}\right\rVert_{L^{\infty}}\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}\left\lVert\eta\right\rVert_{L^{2}}$
    $\leq\frac{\epsilon}{4}\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}+\frac{\left\lVert\mathbf{k}\right\rVert_{L^{\infty}}^{2}}{\epsilon}\left\lVert\eta\right\rVert_{L^{2}}^{2}$
    $\leq\frac{\epsilon}{4}\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}+\frac{\left\lVert\mathbf{k}\right\rVert_{L^{\infty}}^{2}C_{I}^{2}N^{-2}}{\epsilon}\left\lVert v_{N}^{-}\right\rVert_{H^{1}}^{2}.$
Combining these bounds for $|a(v_{N}^{-},\eta)|$:
$|a(v_{N}^{-},\eta)|\leq\left(\frac{\epsilon}{4}+C_{O}N^{-1}+\frac{C^{\prime}N^{-2}}{\epsilon}\right)\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}+\left(C_{O}N^{-1}+\frac{C^{\prime}N^{-2}}{\epsilon}\right)\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}.$
For $N$ large enough such that $C_{O}N^{-1}<\epsilon/4$, the coefficient of $\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}$ is less than $\epsilon/2$. The coefficient of $\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}$ is bounded by a constant.
3. Final Assembly. Summing the bounds for all error components ($E_{1}$ and the parts of $E_{2}$), we find that for any chosen $\epsilon>0$, we can choose $N$ large enough such that the total contribution to the $\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}$ term is less than $\epsilon$. All other terms can be collected into a single term $C_{\epsilon}\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}$, where $C_{\epsilon}$ depends on $\epsilon$, operator norms, and interpolation constants, but is independent of $v_{N}$ and $N$. This completes the proof. ∎