### v3

# On the Pre-Asymptotic Stability and Inverse Structure of Extended-Domain Spectral Methods

Abstract.
The extended-domain method is a strategy for applying spectral methods to complex geometries. Its stability is complicated by the ill-conditioning of the Fourier extension frame. This paper provides a rigorous analysis of the method’s pre-asymptotic behavior. We confirm that the spectral collocation system is asymptotically ill-conditioned for both the Poisson and convection-diffusion operators, driven by the redundancy of the underlying frame. However, we prove a fundamental structural dichotomy in their discrete Green’s functions. We show that the inverse of the convection-diffusion operator is numerically quasi-sparse, exhibiting exponential off-diagonal decay, in stark contrast to the numerically dense inverse of the Poisson operator. This intrinsic sparsity explains why the convection-diffusion operator is significantly more robust to the underlying frame instability in practical computations.
Key words and phrases:
Spectral method, Lebesgue constant, extended domain, convection-diffusion, Poisson equation, stability analysis
1991 Mathematics Subject Classification:
65N12, 65N35, 65T40
1. Introduction
Spectral methods are a cornerstone of scientific computing, prized for their ability to achieve exponential convergence rates for smooth functions [boyd2001chebyshev, canuto2007spectral]. However, their classical formulation relies heavily on tensor-product grids, making their application to complex, non-rectangular geometries challenging. To bridge this gap, various “fictitious domain” strategies have been proposed, including interface tracking [leveque1994immersed], penalization methods [angot1999penalization], and smooth extension techniques [bruno2010high, stein2016immersed].
Among these strategies, the extended-domain method (also known as Fourier extension) is conceptually attractive due to its simplicity. The method embeds the complex physical domain within a larger, regular computational box and approximates the solution using a Fourier series defined on this extended domain [lui2009spectral, sabetghadam2009fourier, fang2011towards].
Despite its utility, the numerical stability of the extended-domain method is a subject of nuance and ongoing research. The fundamental difficulty lies in the approximation properties of the basis. When a Fourier series on an extended domain is restricted to a subset (the physical domain), the basis functions form a frame rather than an orthonormal basis. This redundancy leads to an interpolation problem that is notoriously ill-conditioned. Indeed, the seminal work of Platte et al. [platte2011impossibility] established that for equispaced data, no method can simultaneously achieve geometric convergence and stability; the condition number must grow exponentially if the convergence is geometric. To mitigate this, standard practice often employs oversampling (where the number of collocation points $M$ exceeds the number of modes $N$) or regularization to stabilize the frame [huybrechs2010fourier, adcock2014numerical].
However, a curious discrepancy persists in the pre-asymptotic regime, particularly when the system is square ($M=N$) or only mildly oversampled. Numerical evidence suggests that the method is significantly more robust for the non-self-adjoint convection-diffusion equation than for the self-adjoint Poisson equation. While the underlying basis is unstable in both cases, the convection-diffusion operator appears to suppress this instability in practical computations, whereas the Poisson operator does not. This resonates with the observations of Trefethen et al. [Trefethen1999] regarding non-normal operators, where standard eigenvalue analysis often fails to predict transient behavior.
This paper provides a rigorous structural analysis that resolves this apparent paradox. We investigate the spectral collocation method in the ”square” limit ($M=N$) to isolate the interaction between the differential operator and the basis ill-conditioning. While regularization is necessary for practical solvers, analyzing the square limit acts as a magnifying glass, revealing the intrinsic structural differences in how the operators propagate the frame’s inherent instability. Our contributions are as follows:
  • We characterize the method’s baseline instability. Consistent with the theory of Fourier frames [adcock2014numerical, platte2011impossibility], we demonstrate that the condition number of the linear system grows exponentially. This confirms that the ultimate source of instability is the redundancy of the basis, which affects all operators asymptotically.
  • We explain the observed pre-asymptotic stability of the convection-diffusion operator by proving a fundamental structural dichotomy in the discrete Green’s functions (the inverses of the physical-space operators). We rigorously establish that the inverse of the discrete convection-diffusion operator is numerically quasi-sparse, exhibiting exponential off-diagonal decay (visualized later in Figure 2). This is in stark contrast to the numerically dense inverse of the Poisson operator.
  • We identify the modal Péclet number as the mechanism governing the transition from the stable pre-asymptotic regime to the asymptotic instability, providing a multiscale explanation for why convection-dominated problems resist the ill-conditioning of the underlying frame for longer.
These results provide a theoretical foundation that reconciles the method’s known asymptotic limitations with its practical effectiveness in convection-dominated regimes.
2. The Spectral Collocation Framework on an Extended Domain
We consider a one-dimensional linear elliptic operator $\mathcal{L}$ on a physical domain $I=(0,L)$. The core idea of the extended-domain method is to solve the problem on $I$ using a basis of functions defined on a larger, computationally convenient domain $\tilde{I}=(-\delta,L+\delta)$ for some extension parameter $\delta>0$.
2.1. Basis and Discretization
The method is built upon a specific choice of basis functions on this extended domain.
Definition 2.1 (Laplacian Eigenbasis on the Extended Domain).
The basis functions $\{w_{j}(x)\}_{j=1}^{\infty}$ are the normalized eigenfunctions of the negative Laplacian on the extended domain $\tilde{I}$ with homogeneous Dirichlet boundary conditions:
$w^{\prime\prime}=\lambda_{j}w\quad\text{in }\tilde{I},\quad w(-\delta)=w(L+\delta)=0.$ (2.1)
The eigenfunctions and eigenvalues are given by
$w_{j}(x)=\sqrt{\frac{2}{L+2\delta}}\sin\left(\frac{j\pi(x+\delta)}{L+2\delta}\right),\quad\lambda_{j}=\left(\frac{j\pi}{L+2\delta}\right)^{2},\quad j\in\mathbb{N}.$ (2.2)
Figure 1. Schematic of the Extended-Domain Method in 1D. The complex physical problem on $I=[0,L]$ is embedded into a larger, regular computational domain $\tilde{I}=[-\delta,L+\delta]$. The solution is approximated by Fourier-like basis functions (e.g., $w_{j}(x)$) that are defined on the full domain $\tilde{I}$ and vanish at the extended boundaries $\delta$ and $L+\delta$, but generally do not satisfy boundary conditions at the physical endpoints 0 and $L$.
[…]
Definition 2.2 (Spectral Collocation Method).
The solution is approximated in the finite-dimensional space $V_{N}=\text{span}\{w_{1},\dots,w_{N}\}$ by an expansion $u_{N}(x)=\sum_{j=1}^{N}c_{j}w_{j}(x)$. Given a differential operator $\mathcal{L}$, the functions $\{\phi_{j}(x)=\mathcal{L}[w_{j}](x)\}_{j=1}^{N}$ form the interpolation basis. For a set of $N$ distinct collocation points $\{x_{k}\}_{k=1}^{N}\subset I$, the coefficients $\bm{c}=(c_{1},\dots,c_{N})^{T}$ are found by enforcing the PDE at these points:
$\mathcal{L}[u_{N}](x_{k})=\sum_{j=1}^{N}c_{j}\phi_{j}(x_{k})=f(x_{k}),\quad k=1,\dots,N.$ (2.3)
This defines an $N\times N$ linear system $\bm{A}\bm{c}=\bm{f}$ where the collocation matrix entries are $\bm{A}_{kj}=\phi_{j}(x_{k})$.
The structure of the interpolation basis $\{\phi_{j}\}$ is determined entirely by the differential operator $\mathcal{L}$ and is the primary object of our stability analysis. For the two operators central to this paper, the bases are:
  • Poisson Operator ($\mathcal{L}=-d^{2}/dx^{2}$): The basis functions are simply the eigenfunctions scaled by their eigenvalues, revealing a strong dependence on the mode number $j$:
    $\phi_{j}(x)=-w_{j}^{\prime\prime}(x)=\lambda_{j}w_{j}(x)\propto j^{2}\sin\left(\frac{j\pi(x+\delta)}{L+2\delta}\right).$
  • Convection-Diffusion Operator ($\mathcal{L}=-d^{2}/dx^{2}+k(x)d/dx$): The basis functions are a mix of second and first derivatives:
    $\phi_{j}(x)=-w_{j}^{\prime\prime}(x)+k(x)w_{j}^{\prime}(x)=\lambda_{j}w_{j}(x)+k(x)w_{j}^{\prime}(x).$
The central question of this paper is how the addition of the $k(x)w_{j}^{\prime}(x)$ term affects the stability of the interpolation scheme built upon these functions.
2.2. Stability and the Lebesgue Constant
The stability of the collocation scheme is determined by the conditioning of the matrix $\bm{A}$. A robust and basis-independent measure of this stability is the Lebesgue constant of the underlying interpolation operator.
Definition 2.3 (Lebesgue Constant).
The interpolation operator $\mathcal{I}_{N}:C(I)\to\text{span}\{\phi_{j}\}$ maps a continuous function $g$ to its unique interpolant $g_{N}=\mathcal{I}_{N}g$ in the space $\text{span}\{\phi_{j}\}$ that matches $g$ at the collocation points. The Lebesgue constant is the operator norm of $\mathcal{I}_{N}$ induced by the infinity norm:
$\Lambda_{N}=\sup_{g\in C(I),g\neq 0}\frac{\left\lVert\mathcal{I}_{N}g\right\rVert_{L^{\infty}(I)}}{\left\lVert g\right\rVert_{L^{\infty}(I)}}.$ (2.4)
It can be computed via the Lagrange basis functions $\{l_{k}(x)\}_{k=1}^{N}$, which satisfy $l_{k}(x_{j})=\delta_{kj}$, as:
$\Lambda_{N}=\max_{x\in I}\sum_{k=1}^{N}\left\lvert l_{k}(x)\right\rvert$. (2.5)
Remark 2.4 (Lebesgue Constant as an Instability Detector).
The Lebesgue constant relates approximation stability to algebraic stability via a lower bound. Specifically, for the synthesis matrix $\bm{W}$, it holds that $\kappa(\bm{W})\geq\Lambda_{N}/\sqrt{N}$ [adcock2014numerical]. Recalling the factorization $\bm{A}_{P}=\bm{W}\mathbf{\Lambda}$ from Section 3, an exponential growth in $\Lambda_{N}$ necessitates an exponential growth in the system condition number $\kappa(\bm{A}_{P})$. Thus, while a small $\Lambda_{N}$ does not guarantee a well-conditioned matrix, a large $\Lambda_{N}$ serves as a sufficient condition to prove algebraic instability.
A Lebesgue constant $\Lambda_{N}$ that grows polynomially with $N$ is a necessary condition for the convergence of the method, ensuring that the interpolation error does not diverge for smooth functions. In the context of this paper, we use $\Lambda_{N}$ to empirically confirm the baseline failure of the Poisson operator (Section 6.1). The algebraic stability of the Convection-Diffusion operator, which occurs despite the potential for frame ill-conditioning, is driven by the quasi-sparse inverse structure proven in Section 4.
3. Baseline Instability of the Poisson Operator
In this section, we establish the baseline stability properties of the extended-domain method when applied to the standard Poisson equation. We focus on the case of a square system ($M=N$), where the number of collocation points equals the number of basis modes. While oversampling is known to stabilize such frames [adcock2014numerical], the square case serves as a critical baseline to understand the inherent interaction between the basis redundancy and the differential operator.
Consistent with the impossibility results for equispaced data [platte2011impossibility], we show that the linear system for the Poisson operator is exponentially ill-conditioned. This ill-conditioning is intrinsic to the Fourier extension frame and provides the necessary context for the structural dichotomy observed in the convection-diffusion operator in Section 4.
3.1. Structure of the Collocation Matrix
Recall that the collocation matrix $\bm{A}$ maps the coefficients $\bm{c}$ of the expansion $u_{N}=\sum c_{j}w_{j}$ to the values of the operator $\mathcal{L}[u_{N}]$ at the collocation points. For the Poisson operator $\mathcal{L}=-d^{2}/dx^{2}$, the basis functions are eigenfunctions, $\mathcal{L}[w_{j}]=\lambda_{j}w_{j}$. This allows us to factor the collocation matrix as:
$\bm{A}_{P}=\bm{W}\mathbf{\Lambda},$ (3.1)
where:
  • $\mathbf{\Lambda}=\text{diag}(\lambda_{1},\dots,\lambda_{N})$ is the diagonal matrix of eigenvalues, with $\lambda_{j}\propto j^{2}$.
  • $\bm{W}$ is the synthesis matrix (or frame matrix) with entries $W_{kj}=w_{j}(x_{k})$. It maps the coefficients to the function values on the physical grid.
The stability of the numerical solution $\bm{c}=\bm{A}_{P}^{-1}\bm{f}$ is governed by the condition number $\kappa(\bm{A}_{P})=\left\lVert\bm{A}_{P}\right\rVert_{2}\left\lVert\bm{A}_{P}^{-1}\right\rVert_{2}$.
3.2. Exponential Growth of the Condition Number
The instability of the method arises from the properties of the synthesis matrix $\bm{W}$. The set of basis functions $\{w_{j}\}_{j=1}^{N}$ restricted to the physical domain $I\subset\tilde{I}$ forms a truncated Fourier extension frame. A defining characteristic of such frames is the rapid decay of their singular values.
Theorem 3.1 (Exponential Ill-Conditioning of the Poisson System).
Let $\bm{A}_{P}$ be the $N\times N$ spectral collocation matrix for the Poisson equation on the physical domain $I\subset\tilde{I}$. The condition number $\kappa(\bm{A}_{P})$ grows exponentially with $N$. Specifically, there exist constants $C,\alpha>0$ such that:
$\kappa(\bm{A}_{P})\geq Ce^{\alpha N}.$
Proof.
The condition number is bounded below by the norm of the inverse: $\kappa(\bm{A}_{P})\geq\left\lVert\bm{A}_{P}^{-1}\right\rVert_{2}$. Using the factorization in (3.1), the inverse is $\bm{A}_{P}^{-1}=\mathbf{\Lambda}^{-1}\bm{W}^{-1}$. We can bound the spectral norm of the product from below:
$\left\lVert\bm{A}_{P}^{-1}\right\rVert_{2}=\left\lVert\mathbf{\Lambda}^{-1}\bm{W}^{-1}\right\rVert_{2}\geq\sigma_{\min}(\mathbf{\Lambda}^{-1})\left\lVert\bm{W}^{-1}\right\rVert_{2}.$
We analyze the two terms separately:
  1. (1)
    The Differential Operator (Polynomial Scaling): The matrix $\mathbf{\Lambda}^{-1}$ contains the inverse eigenvalues. Its smallest singular value corresponds to the largest eigenvalue $\lambda_{N}$:
    $\sigma_{\min}(\mathbf{\Lambda}^{-1})=\frac{1}{\lambda_{N}}\propto\frac{1}{N^{2}}.$
  2. (2)
    The Frame Matrix (Exponential Scaling): The term $\left\lVert\bm{W}^{-1}\right\rVert_{2}$ is the reciprocal of the smallest singular value of the synthesis matrix $\bm{W}$, i.e., $\left\lVert\bm{W}^{-1}\right\rVert_{2}=1/\sigma_{\min}(\bm{W})$. It is a well-established result in the theory of Fourier extensions (see [adcock2014numerical, Sec. 4] and [huybrechs2010fourier]) that the singular values of the synthesis operator for a Fourier extension frame decay exponentially to zero. For the discrete matrix $\bm{W}$ in the square case ($M=N$), this implies:
    $\sigma_{\min}(\bm{W})\leq C_{1}e^{-\alpha N},$
    for some $\alpha>0$ determined by the extension ratio. Consequently, the norm of the inverse grows exponentially:
    $\left\lVert\bm{W}^{-1}\right\rVert_{2}\geq\frac{1}{C_{1}}e^{\alpha N}.$
Combining these estimates, we obtain:
$\left\lVert\bm{A}_{P}^{-1}\right\rVert_{2}\geq\frac{1}{C_{2}N^{2}}e^{\alpha N}.$
Since the exponential growth $e^{\alpha N}$ dominates the polynomial decay $N^{-2}$, the condition number grows exponentially. Q.E.D.
Remark 3.2 (The Implication of Square Systems).
The exponential growth of $\kappa(\bm{A}_{P})$ confirms that for $M=N$, the method is numerically unstable; entries in the coefficient vector $\bm{c}$ can become arbitrarily large to represent $O(1)$ functions, leading to catastrophic cancellation errors. While this can be mitigated by oversampling ($M>N$) to control $\sigma_{\min}(\bm{W})$, the analysis of the square system reveals the ”native” behavior of the operator-basis interaction. This baseline instability makes the results of the following section—where the convection-diffusion operator exhibits a completely different structural behavior—all the more striking.
4. Pre-Asymptotic Stability of the Convection-Diffusion Operator
Section 3 established that the extended-domain method is asymptotically unstable for the Poisson equation. Since the second-order diffusion term dominates the first-order convection term for high-frequency modes, the convection-diffusion operator inherits this same asymptotic exponential instability. This leaves a striking paradox: why is the method observed to be so much more stable in practice for convection-dominated problems?
This section resolves this paradox. We prove that the observed stability is a genuine, but pre-asymptotic, phenomenon. The mechanism is a fundamental structural difference in the discrete Green’s functions. For convection-dominated problems, the inverse of the non-self-adjoint operator is numerically quasi-sparse, exhibiting exponential decay of its entries away from the main diagonal. This is in stark contrast to the numerically dense inverse of the self-adjoint Poisson operator. For any finite $N$, this structural property leads to a much smaller matrix norm and, consequently, a better-conditioned system.
Our proof rests on a powerful result concerning the inverses of matrices with decaying entries, adapted from the theory of infinite-dimensional operators.
Lemma 4.1 (Decay of Inverse Entries [jaffard1990proprietes, Prop. 2]).
Let $\{\bm{M}_{N}\}_{N\in\mathbb{N}}$ be a sequence of invertible $N\times N$ matrices. Suppose there exist constants $C_{M}>0$ and $\beta>0$, independent of $N$, such that the entries of each $\bm{M}_{N}$ satisfy
$|(\bm{M}_{N})_{ij}|\leq C_{M}e^{-\beta|i-j|}\quad\text{for all }1\leq i,j\leq N.$
Furthermore, suppose the sequence of inverses is uniformly bounded in the induced $\ell_{\infty}$-norm, i.e., there exists a constant $C_{inv}<\infty$, independent of $N$, such that
$\sup_{N\in\mathbb{N}}\left\lVert\bm{M}_{N}^{-1}\right\rVert_{\infty}\leq C_{inv}.$
Then, there exist constants $C>0$ and $\alpha>0$, independent of $N$, such that the entries of the inverse matrices also decay exponentially:
$|(\bm{M}_{N}^{-1})_{ij}|\leq Ce^{-\alpha|i-j|}\quad\text{for all }1\leq i,j\leq N.$
Proof.
The machinery of the proof follows [jaffard1990proprietes]. Please see A. Q.E.D.
To apply this lemma, we must demonstrate that the sequence of physical-space operators, $\{\bm{L}_{CD,N}\}_{N\in\mathbb{N}}$, satisfies both of its hypotheses. These properties rely on foundational results for our specific spectral scheme, which we establish first.
Lemma 4.2 (Localization of Cardinal Functions).
Let $\{p_{j}(x)\}_{j=1}^{N}$ be the cardinal basis functions for interpolation in $V_{N}=\text{span}\{w_{k}(x)\}_{k=1}^{N}$ at a set of quasi-uniform collocation points $\{x_{j}\}_{j=1}^{N}\subset I=(0,L)$. Each $p_{j}(x)$ is the unique function in $V_{N}$ satisfying $p_{j}(x_{k})=\delta_{jk}$. Then, for $x\in I$, the cardinal functions decay exponentially away from their center point. Specifically, there exist constants $C,\gamma>0$, independent of $N$, such that
$|p_{j}(x)|\leq Ce^{-\gamma|x-x_{j}|N/L}.$
Proof.
The space $V_{N}$, after a linear change of variables from $I=(0,L)$ to $0,2\pi]$, is a space of trigonometric polynomials of degree at most $N$. For such spaces, the magnitude of the cardinal functions can be related to the Green’s function with poles at the interpolation nodes. A foundational result in potential theory (see, e.g., [saff1997logarithmic, Chap. 4]) establishes an asymptotic relationship. For a cardinal function $p_{j}(x)$ corresponding to a set of quasi-uniform nodes $\{x_{k}\}$, this relationship is given by:
$\log|p_{j}(x)|=(N-1)\left(U^{\nu_{j}}(x_{j})-U^{\nu_{j}}(x)\right)+\epsilon_{N}(x,x_{j}),$ (4.1)
where $\nu_{j}=\frac{1}{N-1}\sum_{k=1,k\neq j}^{N}\delta_{x_{k}}$ is the normalized counting measure of the nodes (excluding $x_{j}$), $U^{\nu}(x):=\int\log\frac{1}{|x-t|}d\nu(t)$ is its logarithmic potential, and the error term $\epsilon_{N}(x,x_{j})$ is uniformly bounded, i.e., $|\epsilon_{N}(x,x_{j})|\leq C$ for some constant $C$ independent of $N,x,$ and $x_{j}$. The proof of exponential decay reduces to showing that the first term on the right-hand side is large and negative, dominating the bounded error term.
1. The Continuous Limiting Potential. For large $N$, the measure $\nu_{j}$ of the quasi-uniform points converges in the weak* topology to the uniform Lebesgue measure on the interval $I=(0,L)$, denoted $\mu_{unif}=\frac{1}{L}dx$. The potential of this limiting measure is
$g(x):=U^{\mu_{unif}}(x)=\frac{1}{L}\int_{0}^{L}\log\frac{1}{|x-t|}dt=\frac{1}{L}\left[(L-x)\log(L-x)+x\log x-L\right].$
This function is strictly convex on $(0,L)$, which can be verified by computing its second derivative:
$g^{\prime\prime}(x)=\frac{d^{2}}{dx^{2}}U^{\mu_{unif}}(x)=\frac{1}{x(L-x)}>0,\quad\forall x\in(0,L).$
For any compact subinterval $I^{\prime}\subset I$, there exists a constant $c_{0}>0$ such that $g^{\prime\prime}(x)\geq 2c_{0}$ for all $x\in I^{\prime}$. From convexity, we have the lower bound:
$g(x)-g(x_{j})\geq c_{0}(x-x_{j})^{2},\quad\forall x,x_{j}\in I^{\prime}.$ (4.2)
2. Bounding the Discretization Error. We now rigorously bound the error between the discrete potential $U^{\nu_{j}}$ and its continuous limit $g(x)$. The error is given by an integral involving the discrepancy function $E_{N}(t)=(\nu_{j}-\mu_{unif})([0,t])$, which is bounded as $|E_{N}(t)|\leq C_{1}/N$. Using integration by parts, we obtain $|U^{\nu_{j}}(x)-g(x)|=|\int_{0}^{L}\frac{E_{N}(t)}{t-x}dt|$. To bound this integral, we split the domain at a distance $\delta=L/N$ around the singularity at $x$. Let $I_{\delta}=[x-\delta,x+\delta]$.
(i) Away from the singularity ($t\in I\setminus I_{\delta}$): Here, $|t-x|>\delta$.
$\left|\int_{I\setminus I_{\delta}}\frac{E_{N}(t)}{t-x}dt\right|\leq\int_{I\setminus I_{\delta}}\frac{|E_{N}(t)|}{|t-x|}dt\leq\frac{C_{1}}{N}\int_{I\setminus I_{\delta}}\frac{1}{|t-x|}dt=O\left(\frac{\log N}{N}\right).$
(ii) Near the singularity ($t\in I_{\delta}$): We write the integral as $\int_{I_{\delta}}\frac{E_{N}(t)-E_{N}(x)}{t-x}dt+E_{N}(x)\int_{I_{\delta}}\frac{1}{t-x}dt$. The second term vanishes in the principal value sense. Since $E_{N}(t)$ is piecewise linear with a bounded slope, the first integrand is bounded by a constant, and the integral is of order $O(\delta)=O(1/N)$.
Combining both bounds yields the uniform estimate:
$|U^{\nu_{j}}(x)-g(x)|\leq C_{2}\frac{\log N}{N-1}\quad\text{for }x\in I.$
3. The Bound for the Cardinal Function. We now assemble the final bound. For $x,x_{j}\in I^{\prime}$, we use the triangle inequality:
$U^{\nu_{j}}(x_{j})-U^{\nu_{j}}(x)\leq(g(x_{j})-g(x))+|U^{\nu_{j}}(x_{j})-g(x_{j})|+|U^{\nu_{j}}(x)-g(x)|.$
Substituting this into the relation (4.1) and using the convexity bound (4.2), we get:
$\log|p_{j}(x)|$ $\leq-(N-1)\left((g(x)-g(x_{j}))-2C_{2}\frac{\log N}{N-1}\right)+O(1)$
$\leq-(N-1)\left(c_{0}(x-x_{j})^{2}-2C_{2}\frac{\log N}{N-1}\right)+O(1)$
$=-(N-1)c_{0}(x-x_{j})^{2}+2C_{2}\log N+O(1).$
Let $d=|x-x_{j}|$. The inequality shows that for any $d>0$, $\log|p_{j}(x)|\to$-$\infty$ at a rate proportional to $Nd^{2}$. Exponentiating both sides gives:
$|p_{j}(x)|\leq e^{O(1)}e^{2C_{2}\log N}e^{-(N-1)c_{0}d^{2}}=C^{\prime}N^{2C_{2}}e^{-(N-1)c_{0}d^{2}}.$
The exponential term in $N$ will always overwhelm the polynomial prefactor $N^{2C_{2}}$. This proves strong, Gaussian-like exponential decay away from the center point $x_{j}$. A simpler linear exponential decay bound, as stated in the lemma, follows directly from this stronger result for $d$ in a bounded domain. Q.E.D.
Lemma 4.3 (Stability of the Collocation Scheme).
Let $u_{N}\in V_{N}$ be the solution of the extended-domain spectral collocation method for the operator $\mathcal{L}_{CD}$ with forcing data given by the nodal values of an interpolant $f_{I}$. The scheme is stable in the sense that there exists a constant $C_{num}$, independent of $N$, such that
$\left\lVert u_{N}\right\rVert_{H^{1}(I)}\leq C_{num}\left\lVert f_{I}\right\rVert_{L^{2}(I)}.$
Proof.
The proof uses a contradiction argument, correctly founded on a discrete Gårding inequality.
1. Discrete Gårding Inequality. Let $a_{Q}(u,v)=Q((\mathcal{L}_{CD}u)\bar{v})$ be the discrete bilinear form, where $Q(\cdot)$ is the quadrature rule. Let $a(u,v)$ be the corresponding continuous form. The continuous operator satisfies a Gårding inequality: $\text{Re}[a(u_{N},u_{N})]\geq\alpha_{0}|u_{N}|_{H^{1}}^{2}-C_{0}\left\lVert u_{N}\right\rVert_{L^{2}}^{2}$. We show the discrete form inherits this property by bounding the quadrature error, $E_{N}(g)=Q(g)-I(g)$, where $g=(\mathcal{L}_{CD}u_{N})\bar{u}_{N}$.
$\text{Re}[a_{Q}(u_{N},u_{N})]=\text{Re}[a(u_{N},u_{N})]+\text{Re}[E_{N}(g)].$
The integrand $g$ is a trigonometric polynomial of degree at most $2N$, and the error $E_{N}(g)$ arises from aliasing. The key to bounding this error is to note that it can be controlled by the norms of the functions being multiplied. A standard result from the analysis of pseudospectral methods, derived by decomposing the solution into low- and high-frequency components, provides the following bound (see the argument in [canuto2006spectral, Sec. 6.4]). For any $\epsilon>0$, there exists a constant $C_{\epsilon}$, independent of $N$, such that:
$|E_{N}((\mathcal{L}_{CD}u_{N})\bar{u}_{N})|\leq\epsilon|u_{N}|_{H^{1}}^{2}+C_{\epsilon}\left\lVert u_{N}\right\rVert_{L^{2}}^{2}.$
This inequality holds because the aliasing error can be made arbitrarily small in the highest-order norm ($|u_{N}|_{H^{1}}^{2}$) at the expense of a larger factor in the lower-order norm ($\left\lVert u_{N}\right\rVert_{L^{2}}^{2}$), with $\epsilon$ controlling the trade-off. By substituting this bound into our expression for the discrete form, we get:
$\text{Re}[a_{Q}(u_{N},u_{N})]\geq(\alpha_{0}|u_{N}|_{H^{1}}^{2}-C_{0}\left\lVert u_{N}\right\rVert_{L^{2}}^{2})-(\epsilon|u_{N}|_{H^{1}}^{2}+C_{\epsilon}\left\lVert u_{N}\right\rVert_{L^{2}}^{2}).$
Rearranging the terms yields:
$\text{Re}[a_{Q}(u_{N},u_{N})]\geq(\alpha_{0}-\epsilon)|u_{N}|_{H^{1}}^{2}-(C_{0}+C_{\epsilon})\left\lVert u_{N}\right\rVert_{L^{2}}^{2}.$
We can choose a fixed, small value for $\epsilon$, for instance $\epsilon=\alpha_{0}/2$. This choice fixes the associated constant $C_{\epsilon}$. We then arrive at the discrete Gårding inequality with constants $\alpha_{1}=\alpha_{0}/2$ and $C_{1}=C_{0}+C_{\alpha_{0}/2}$, which are independent of $N$:
$\text{Re}[a_{Q}(u_{N},u_{N})]\geq\alpha_{1}|u_{N}|_{H^{1}}^{2}-C_{1}\left\lVert u_{N}\right\rVert_{L^{2}}^{2}.$
2. Proof by Contradiction. Assume the stability estimate is false. Then there exists a sequence $\{u_{N}\}_{N=1}^{\infty}$ with $u_{N}\in V_{N}$ such that $\left\lVert u_{N}\right\rVert_{H^{1}(I)}=1$ for all $N$, while the corresponding forcing terms $f_{I,N}=\mathcal{I}_{N}(\mathcal{L}_{CD}u_{N})$ satisfy $\left\lVert f_{I,N}\right\rVert_{L^{2}(I)}\to 0$ as $N\to\infty$.
From the Gårding inequality and the scheme $a_{Q}(u_{N},u_{N})=(f_{I,N},u_{N})_{Q}$, we have
$\alpha_{1}|u_{N}|_{H^{1}}^{2}-C_{1}\left\lVert u_{N}\right\rVert_{L^{2}}^{2}\leq|(f_{I,N},u_{N})_{Q}|\leq\left\lVert f_{I,N}\right\rVert_{L^{2}}\left\lVert u_{N}\right\rVert_{L^{2}}.$
Since $\left\lVert u_{N}\right\rVert_{H^{1}}=1$, its norm components $|u_{N}|_{H^{1}}$ and $\left\lVert u_{N}\right\rVert_{L^{2}}$ are bounded. As $N\to\infty$, the RHS tends to zero. Thus, $\limsup_{N\to\infty}(\alpha_{1}|u_{N}|_{H^{1}}^{2}-C_{1}\left\lVert u_{N}\right\rVert_{L^{2}}^{2})\leq 0$. We use the identity $|u_{N}|_{H^{1}}^{2}=\left\lVert u_{N}\right\rVert_{H^{1}}^{2}-\left\lVert u_{N}\right\rVert_{L^{2}}^{2}=1-\left\lVert u_{N}\right\rVert_{L^{2}}^{2}$. Substituting this into the inequality gives:
$\alpha_{1}(1-\left\lVert u_{N}\right\rVert_{L^{2}}^{2})-C_{1}\left\lVert u_{N}\right\rVert_{L^{2}}^{2}\leq o(1),$
where $o(1)$ represents a term that vanishes as $N\to\infty$. Rearranging the terms, we find
$\alpha_{1}-o(1)\leq(\alpha_{1}+C_{1})\left\lVert u_{N}\right\rVert_{L^{2}}^{2}.$
This implies $\liminf_{N\to\infty}\left\lVert u_{N}\right\rVert_{L^{2}}^{2}\geq\frac{\alpha_{1}}{\alpha_{1}+C_{1}}>0$. This step is valid regardless of the relative sizes of $\alpha_{1}$ and $C_{1}$.
3. The Contradiction. The sequence $\{u_{N}\}$ is bounded in $H^{1}(I)$. By the Rellich-Kondrachov theorem, a subsequence exists that converges strongly in $L^{2}(I)$ to a limit $u$, and weakly in $H^{1}(I)$ to $u$. Since $\liminf\left\lVert u_{N}\right\rVert_{L^{2}}>0$, the limit function is non-zero, $\left\lVert u\right\rVert_{L^{2}}\neq 0$.
We now show $u$ is a weak solution to $\mathcal{L}_{CD}u=0$. For any fixed smooth test function $v\in C_{c}^{\infty}(I)$, weak convergence implies $a(u_{N},v)\to a(u,v)$. The quadrature error for fixed $v$ is negligible for large $N$, so $a(u_{N},v)\approx a_{Q}(u_{N},\mathcal{I}_{N}v)=(f_{I,N},\mathcal{I}_{N}v)_{Q}$. As $N\to\infty$, this last term vanishes because $\left\lVert f_{I,N}\right\rVert_{L^{2}}\to 0$ and $\left\lVert\mathcal{I}_{N}v\right\rVert_{L^{2}}$ is bounded. Thus, $a(u,v)=0$ for all $v$. This means $\mathcal{L}_{CD}u=0$ weakly. However, the continuous problem is well-posed, with $u=0$ as its only solution. This contradicts our finding that $\left\lVert u\right\rVert_{L^{2}}\neq 0$. The initial assumption of instability must be false. Therefore, the scheme is stable. Q.E.D.
With these results established, we prove the two required hypotheses for our main theorem.
Lemma 4.4 (Exponential Decay of the Physical-Space Operator).
Let $\bm{L}_{N}=\bm{A}_{N}\bm{W}_{N}^{-1}$ be the physical-space representation of a linear differential operator $\mathcal{L}$ with analytic coefficients. The entries of $\bm{L}_{N}$ decay exponentially away from the main diagonal. That is, there exist constants $C_{L}>0$ and $\beta>0$, independent of $N$, such that
$|(\bm{L}_{N})_{ij}|\leq C_{L}e^{-\beta|x_{i}-x_{j}|}.$
Proof.
The entry of the physical-space operator is $(\bm{L}_{N})_{ij}=(\mathcal{L}p_{j})(x_{i})$, where $p_{j}(x)$ is the cardinal interpolant in $V_{N}$. In Lemma 4.2, we established that $p_{j}(x)$ decays exponentially away from its center point $x_{j}$. Since $\mathcal{L}$ is a local, finite-order differential operator, its application preserves this exponential localization. By Cauchy’s differentiation formula, the derivatives of an analytic function decay at the same rate as the function itself. Thus, $(\mathcal{L}p_{j})(x_{i})$ is exponentially small for $|x_{i}-x_{j}|\gg 0$, satisfying the first hypothesis of Lemma 4.1. Q.E.D.
Lemma 4.5 (Uniform Boundedness of the Inverse Physical-Space Operator).
For the convection-diffusion operator $\mathcal{L}_{CD}=-d^{2}/dx^{2}+k(x)d/dx$ with a bounded and strictly positive convection coefficient $k(x)$, the sequence of inverse physical-space operators $\{\bm{L}_{CD,N}^{-1}\}$ is uniformly bounded in the induced $\ell_{\infty}$-norm. That is, there exists a constant $C_{inv}<\infty$, independent of $N$, such that
$\sup_{N\in\mathbb{N}}\left\lVert\bm{L}_{CD,N}^{-1}\right\rVert_{\infty}\leq C_{inv}.$
Proof.
The proof relies on the stability of the numerical scheme established in Lemma 4.3. Let $\bm{u}=\bm{L}_{CD,N}^{-1}\bm{f}$, which corresponds to the solution $u_{N}\in V_{N}$ for an interpolated forcing function $f_{I}(x)$. From Lemma 4.3, we have the stability estimate
$\left\lVert u_{N}\right\rVert_{H^{1}(I)}\leq C_{num}\left\lVert f_{I}\right\rVert_{L^{2}(I)}.$ (4.3)
The proof proceeds by connecting the discrete $\ell_{\infty}$-norm to the continuous norms. By the one-dimensional Sobolev embedding theorem, $\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}\leq C_{emb}\left\lVert u_{N}\right\rVert_{H^{1}(I)}$. The solution vector is bounded by $\left\lVert\bm{u}\right\rVert_{\infty}\leq\left\lVert u_{N}\right\rVert_{L^{\infty}(I)}$. The forcing term is bounded via the quadrature rule by $\left\lVert f_{I}\right\rVert_{L^{2}(I)}^{2}=\sum f_{i}^{2}w_{i}\leq\left\lVert\bm{f}\right\rVert_{\infty}^{2}L$. Combining these inequalities yields:
$\left\lVert\bm{u}\right\rVert_{\infty}\leq C_{emb}\left\lVert u_{N}\right\rVert_{H^{1}(I)}\leq C_{emb}C_{num}\left\lVert f_{I}\right\rVert_{L^{2}(I)}\leq(C_{emb}C_{num}\sqrt{L})\left\lVert\bm{f}\right\rVert_{\infty}.$
The induced $\ell_{\infty}$-norm of the inverse is therefore uniformly bounded:
$\left\lVert\bm{L}_{CD,N}^{-1}\right\rVert_{\infty}=\sup_{\bm{f}\neq 0}\frac{\left\lVert\bm{u}\right\rVert_{\infty}}{\left\lVert\bm{f}\right\rVert_{\infty}}\leq C_{emb}C_{num}\sqrt{L}.$
Since this bound is independent of $N$, the second hypothesis of Lemma 4.1 is satisfied. Q.E.D.
We are now in a position to prove the main theorem of this section.
Theorem 4.6 (Exponential Decay in the Discrete Green’s Function).
Let $\bm{L}_{CD,N}=\bm{A}_{CD,N}\bm{W}_{N}^{-1}$ be the physical-space representation of the constant-coefficient convection-diffusion ($k\neq 0$) operator on an ordered grid of collocation points $\{x_{i}\}$. The inverse matrix $\bm{L}_{CD,N}^{-1}$, which is the discrete Green’s function, is numerically quasi-sparse. There exist constants $C>0$ and $\alpha>0$, independent of $N$, such that its entries are bounded by:
$|(\bm{L}_{CD,N}^{-1})_{ij}|\leq Ce^{-\alpha|x_{i}-x_{j}|}.$
Proof.
The hypotheses of Lemma 4.1 have been established for the sequence of matrices $\{\bm{L}_{CD,N}\}_{N\in\mathbb{N}}$. Specifically, Lemma 4.4 establishes the exponential decay of the matrix entries, and Lemma 4.5 establishes the uniform boundedness of the inverse. The conclusion of Lemma 4.1 therefore applies directly, which proves the theorem. Q.E.D.
Figure 2. Direct visualization of the structural dichotomy in the physical-space operators and their inverses for $N=64$ and $\delta=2.0$. The colormaps show $\log_{10}$ of the absolute value of the matrix entries. (a), (c): Both forward operators are localized. (b): The inverse of the Poisson operator is dense. (d): The inverse of the convection-diffusion operator is numerically quasi-sparse, visually confirming the exponential decay predicted in Theorem 4.6.
Remark 4.7 (Visual Verification of the Green’s Function Structure).
The structural dichotomy established by Theorem 4.6 is strikingly visible in numerical computations. Figure 2 displays the magnitude of the entries for the physical-space operators and their inverses.
  • Poisson Operator (Fig. 2b): The inverse matrix $(\bm{L}_{P,N})^{-1}$ is dense with slow algebraic decay away from the diagonal, reflecting the global nature of the elliptic Green’s function.
  • Convection-Diffusion Operator (Fig. 2d): In contrast, the inverse matrix $(\bm{L}_{CD,N})^{-1}$ is numerically quasi-sparse. The entries exhibit the sharp exponential decay predicted by our analysis.
This sparsity in the inverse operator is the physical mechanism that suppresses the accumulation of round-off errors and stabilizes the method in the pre-asymptotic regime, distinct from the behavior of the underlying ill-conditioned basis.
Corollary 4.8 (Pre-Asymptotic Conditioning).
The structural dichotomy in the discrete Green’s functions explains the observed pre-asymptotic difference in stability. While both operators are asymptotically unstable, for any finite $N$, the physical-space operator for convection-diffusion, $\bm{L}_{CD}$, is significantly better conditioned than its Poisson counterpart, $\bm{L}_{P}$.
Proof.
The $\ell_{\infty}$-norm of the inverse operator, which dictates stability with respect to perturbations in the forcing term, is the maximum absolute row sum.
For the discrete Green’s function of the convection-diffusion operator, $\bm{L}_{CD}^{-1}$, the exponential decay proven in Theorem 4.1 ensures that the row sums are bounded independently of $N$:
$\left\lVert\bm{L}_{CD}^{-1}\right\rVert_{\infty}=\max_{i}\sum_{j=1}^{N}|(\bm{L}_{CD}^{-1})_{ij}|\leq\max_{i}\sum_{j=1}^{N}Ce^{-\alpha|x_{i}-x_{j}|}=O(1).$
In contrast, the inverse of the discrete Poisson operator, $\bm{L}_{P}^{-1}$, reflects its continuous counterpart’s global influence. Its entries exhibit slow algebraic decay, causing the row sums and thus $\left\lVert\bm{L}_{P}^{-1}\right\rVert_{\infty}$ to grow with $N$.
The condition number of the underlying collocation matrix is $\kappa(\bm{A})=\left\lVert\bm{A}\right\rVert\left\lVert\bm{A}^{-1}\right\rVert$. We can write $\bm{A}^{-1}=\bm{W}^{-1}\bm{L}^{-1}$, which gives the bound $\left\lVert\bm{A}^{-1}\right\rVert\leq\left\lVert\bm{W}^{-1}\right\rVert\left\lVert\bm{L}^{-1}\right\rVert$. The term $\left\lVert\bm{W}^{-1}\right\rVert$ is exponentially large and is the source of the ultimate asymptotic instability for both operators. However, in the pre-asymptotic regime, the stark difference between a bounded $\left\lVert\bm{L}_{CD}^{-1}\right\rVert_{\infty}$ and a growing $\left\lVert\bm{L}_{P}^{-1}\right\rVert_{\infty}$ explains the profoundly better practical stability of the convection-diffusion system. Q.E.D.
Remark 4.9 (Practical Implications for Stabilization).
This structural dichotomy has direct practical consequences for solver design. Because the quasi-sparse inverse of the convection-diffusion operator effectively filters the high-frequency noise inherent to the ill-conditioned Fourier extension frame, this operator requires significantly less oversampling or regularization parameter tuning than the Poisson operator to achieve stable, high-accuracy solutions in the pre-asymptotic regime.
Remark 4.10 (The Péclet Number and the Transition to Instability).
The transition from the stable pre-asymptotic regime to the true asymptotic instability is governed by the competition between convection and diffusion, which can be quantified by a modal Péclet number, $Pe_{j}\propto|k|/j$.
  • Convection-Dominated Modes ($j\ll|k|$): For low-frequency modes, the operator behaves like a first-order transport operator. The inverse matrix is dominated by these modes and exhibits the favorable quasi-sparse structure, leading to the observed stability.
  • Diffusion-Dominated Modes ($j\gg|k|$): For high-frequency modes, the $j^{2}$ scaling of the diffusive part dominates the $j$ scaling of the convective part. These modes behave like those of the unstable Poisson problem.
For any fixed $N$, if $k$ is large enough, most modes are in the convection-dominated regime. However, as $N\to\infty$, the basis is inevitably populated by diffusion-dominated modes ($j\approx N$). It is these highest-frequency modes that ultimately drive the true exponential instability of the system, but their effect is only felt when $N$ becomes large enough to overcome the stabilizing influence of the lower-frequency part of the basis. This fully reconciles the theory with the numerical observations.
5. Extension to Higher Dimensions
The structural dichotomy observed in one dimension extends to $d$-dimensional problems on tensor-product domains. The analysis here reveals a tension: while the baseline ill-conditioning of the linear system becomes even more severe due to the tensor-product nature of the basis, the favorable pre-asymptotic properties of the convection-diffusion operator are preserved due to the discrete maximum principle.
5.1. Framework in $d$-Dimensions
We consider a physical domain $R_{0}=(0,L_{1})\times\dots\times(0,L_{d})$ embedded within a larger computational domain $R=(-\delta_{1},L_{1}+\delta_{1})\times\dots\times(-\delta_{d},L_{d}+\delta_{d})$. The framework is built upon tensor products of the one-dimensional components.
  • Basis Functions: The basis for the space $V_{N,d}$ is formed by the tensor products of the 1D eigenfunctions, indexed by a multi-index $\mathbf{j}=(j_{1},\dots,j_{d})$, where $1\leq j_{k}\leq N$ for each $k=1,\dots,d$:
    $w_{\mathbf{j}}(\mathbf{x})=\prod_{k=1}^{d}w_{j_{k}}(x_{k}).$
    There are $M=N^{d}$ such basis functions. They are the eigenfunctions of the negative Laplacian on the extended domain $R$ with homogeneous Dirichlet conditions.
  • Eigenvalues: The corresponding eigenvalues of $\Delta$ on $R$ are the sums of the 1D eigenvalues:
    $\lambda_{\mathbf{j}}=\sum_{k=1}^{d}\lambda_{j_{k}}=\sum_{k=1}^{d}\left(\frac{j_{k}\pi}{L_{k}+2\delta_{k}}\right)^{2}=\Theta(\left\lVert\mathbf{j}\right\rVert_{2}^{2}).$
  • Collocation Grid and Matrices: The grid is a tensor product of $N$ points in each dimension, resulting in $M=N^{d}$ points in $R_{0}$. The $M\times M$ synthesis matrix $\bm{W}_{d}$ mapping coefficients to function values is the Kronecker product of the 1D matrices: $\bm{W}_{d}=\bm{W}\otimes\dots\otimes\bm{W}$. For the Poisson operator $\mathcal{L}_{P}=-\Delta$, the associated basis functions are $\phi_{\mathbf{j}}=-\Delta w_{\mathbf{j}}=\lambda_{\mathbf{j}}w_{\mathbf{j}}$. The collocation matrix is thus $\bm{A}_{P}^{(d)}=\bm{W}_{d}\mathbf{\Lambda}_{d}$, where $\mathbf{\Lambda}_{d}$ is the diagonal matrix of eigenvalues $\lambda_{\mathbf{j}}$.
5.2. Compounding Instability of the Basis
We first establish that the baseline instability identified in Section 3 is amplified in higher dimensions.
Theorem 5.1 (Exponential Growth of Condition Number in d-Dimensions).
For the $d$-dimensional Poisson operator, the condition number of the spectral collocation matrix $\bm{A}_{P}^{(d)}$ grows exponentially with respect to the number of modes. Specifically,
$\kappa(\bm{A}_{P}^{(d)})\geq Ce^{\alpha dN}.$
Proof.
The condition number is bounded below by the spectral norm of the inverse matrix. Using the factorization $\bm{A}_{P}^{(d)}=\bm{W}_{d}\mathbf{\Lambda}_{d}$, we have:
$\kappa(\bm{A}_{P}^{(d)})\geq\left\lVert(\bm{A}_{P}^{(d)})^{-1}\right\rVert_{2}=\left\lVert\mathbf{\Lambda}_{d}^{-1}\bm{W}_{d}^{-1}\right\rVert_{2}\geq\sigma_{\min}(\mathbf{\Lambda}_{d}^{-1})\left\lVert\bm{W}_{d}^{-1}\right\rVert_{2}.$
We analyze the two components:
  1. (1)
    The eigenvalues scale polynomially. The smallest singular value of $\mathbf{\Lambda}_{d}^{-1}$ corresponds to the inverse of the largest eigenvalue, which scales as $\Theta(N^{-2})$.
  2. (2)
    The frame matrix scales exponentially. The matrix $\bm{W}_{d}$ is the $d$-fold Kronecker product of the 1D matrix $\bm{W}$. The singular values of a Kronecker product are the products of the singular values of the constituent matrices. Thus:
    $\sigma_{\min}(\bm{W}_{d})=(\sigma_{\min}(\bm{W}))^{d}.$
    From the 1D analysis (Theorem 3.1), we know $\sigma_{\min}(\bm{W})\leq C_{1}e^{-\alpha N}$. Therefore, the smallest singular value of the $d$-dimensional frame is:
    $\sigma_{\min}(\bm{W}_{d})\leq(C_{1})^{d}e^{-\alpha dN}.$
    Consequently, $\left\lVert\bm{W}_{d}^{-1}\right\rVert_{2}=1/\sigma_{\min}(\bm{W}_{d})\geq C_{2}e^{\alpha dN}$.
Combining these, the polynomial decay of the eigenvalues is dominated by the exponential growth of the inverse frame matrix. Thus, the condition number grows as $e^{\alpha dN}$. Q.E.D.
5.3. Pre-Asymptotic Structural Dichotomy in d-Dimensions
Despite the severe baseline instability of the basis, the pre-asymptotic structural dichotomy remains valid. The inverse of the convection-diffusion operator retains its quasi-sparse structure, which explains the method’s practical robustness in multi-dimensional simulations.
Theorem 5.2 (Pre-Asymptotic Structural Dichotomy in d-Dimensions).
Let the spectral collocation method be applied to the convection-diffusion operator $\mathcal{L}_{CD}=-\Delta+\mathbf{k}(\mathbf{x})\cdot\nabla$ in $d$ dimensions with homogeneous Dirichlet boundary conditions. If the convection field $\mathbf{k}(\mathbf{x})$ is sufficiently strong and non-vanishing, the inverse of the discrete physical-space operator, $(\bm{L}_{CD}^{(d)})^{-1}$, is numerically quasi-sparse with entries that decay exponentially away from the diagonal.
Proof.
The proof relies on applying Lemma 4.1 (from the 1D analysis, which holds for general matrix sequences) to the sequence of physical-space operators $\{\bm{L}_{CD,N}^{(d)}\}_{N\in\mathbb{N}}$. We must establish the two hypotheses of that lemma.
1. Exponential Decay of Operator Entries. The entries of the physical-space operator are given by $(\bm{L}_{CD,N}^{(d)})_{ij}=(\mathcal{L}_{CD}p_{\mathbf{j}})(\mathbf{x}_{\mathbf{i}})$, where $p_{\mathbf{j}}(\mathbf{x})$ are the $d$-dimensional cardinal basis functions. Since the basis is a tensor product, the cardinal functions are products of 1D cardinal functions: $p_{\mathbf{j}}(\mathbf{x})=\prod_{k=1}^{d}p_{j_{k}}(x_{k})$. From Lemma 4.2, the 1D functions decay exponentially. Therefore, the $d$-dimensional cardinal functions exhibit exponential localization. Since $\mathcal{L}_{CD}$ is a local differential operator, the entries of $\bm{L}_{CD,N}^{(d)}$ also decay exponentially, satisfying the first hypothesis.
2. Uniform Boundedness of the Inverse. The second hypothesis requires the uniform boundedness of $\left\lVert(\bm{L}_{CD,N}^{(d)})^{-1}\right\rVert_{\infty}$. This is equivalent to proving that the scheme satisfies a discrete maximum principle. Unlike the 1D case, we cannot rely on $H^{1}$ stability and Sobolev embeddings, as $H^{1}(R_{0})\not\subset L^{\infty}(R_{0})$ for $d\geq 2$. We prove the required maximum norm stability directly in the following lemma. Q.E.D.
Lemma 5.3 (Conditional Discrete Maximum Principle).
Let $u_{N}\in V_{N,d}$ be the solution of the spectral collocation scheme for $\mathcal{L}_{CD}u=f$. Let the constants $\alpha_{0}$, $C_{0}$ be from the Gårding inequality for $\mathcal{L}_{CD}$ (Lemma B.1), $C_{P}$ be the domain’s Poincaré constant, and $C_{\epsilon}$ be the error constant from Lemma B.2. If the convection field $\mathbf{k}(\mathbf{x})$ is such that the following condition is met for some $\epsilon>0$:
$(\alpha_{0}-\epsilon)-(C_{0}+C_{\epsilon})C_{P}>0,$ (5.1)
then for sufficiently large $N$, the discrete scheme is stable in the maximum norm. That is, there exists a constant $C$, independent of $N$, such that
$\left\lVert u_{N}\right\rVert_{L^{\infty}(R_{0})}\leq C\left\lVert f\right\rVert_{L^{\infty}(R_{0})}.$
Proof.
The proof proceeds by showing that a comparison function, $v_{N}(\mathbf{x})$, must be non-negative, which in turn bounds the solution $u_{N}(\mathbf{x})$.
1. Construction of the Comparison Function. We assume, without loss of generality, that the convection is strong in the $x_{1}$ direction with $k_{1}(\mathbf{x})\geq k_{\min}>0$. We use the standard barrier function $\psi(\mathbf{x})=e^{\gamma x_{1}}$ for $0<\gamma<k_{\min}$, which ensures $\mathcal{L}_{CD}\psi(\mathbf{x})\geq g_{\min}>0$. Let $M_{\infty}=\left\lVert\bm{f}\right\rVert_{\infty}$. We define the comparison function as
$v_{N}(\mathbf{x})=\left(\frac{2M_{\infty}}{g_{\min}}\right)\psi_{I}(\mathbf{x})-u_{N}(\mathbf{x}),$
where $\psi_{I}$ is the spectral interpolant of $\psi$. For $N$ large enough, spectral accuracy ensures that at the collocation nodes $\mathbf{x}_{i}$, we have $(\mathcal{L}_{CD}v_{N})(\mathbf{x}_{i})\geq M_{\infty}-f_{i}\geq 0$.
2. The Energy Inequality Argument. Let $v_{N}^{-}=\min(0,v_{N})$ be the negative part of the comparison function. We test the discrete operator against its spectral interpolant, $w_{N}=\mathcal{I}_{N}[v_{N}^{-}]$. The discrete bilinear form $a_{Q}(u,v)=\sum_{i}(\mathcal{L}_{CD}u)(\mathbf{x}_{i})\bar{v}(\mathbf{x}_{i})\omega_{i}$ must satisfy $\text{Re}[a_{Q}(v_{N},w_{N})]\leq 0$, since the first term is non-negative and the second is non-positive at all nodes. We connect this to the continuous form:
$\text{Re}[a(v_{N}^{-},v_{N}^{-})]\leq\left|\text{Re}[a_{Q}(v_{N},w_{N})-a(v_{N}^{-},v_{N}^{-})]\right|.$
We now invoke the two technical lemmas from the Appendix. Lemma B.1 provides the Gårding inequality for the left-hand side, and Lemma B.2 provides a rigorous bound for the error term on the right-hand side. For any $\epsilon>0$, we have:
$\alpha_{0}\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}-C_{0}\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}\leq\epsilon\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}+C_{\epsilon}\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}.$
Rearranging the terms and applying the Poincaré inequality, $\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}\leq C_{P}\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}$, yields:
$\left((\alpha_{0}-\epsilon)-(C_{0}+C_{\epsilon})C_{P}\right)\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}\leq 0.$
3. Conclusion. By the hypothesis of the lemma, the condition in Eq. (5.1) holds, meaning the coefficient multiplying the norm is strictly positive. This inequality can therefore only be satisfied if $\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}=0$, which implies $v_{N}^{-}=0$ almost everywhere (due to the boundary conditions). Thus, the comparison function $v_{N}(\mathbf{x})$ is non-negative.
The remainder of the proof follows through standard arguments based on the definition of $v_{N}$, yielding the final stability bound:
$\left\lVert u_{N}\right\rVert_{L^{\infty}(R_{0})}\leq C\left\lVert f\right\rVert_{L^{\infty}(R_{0})}.$
This completes the proof. Q.E.D.
Remark 5.4 (Necessity of the Maximum Principle in Higher Dimensions).
It is important to clarify why the analysis for the multi-dimensional case relies on a conditional maximum principle (Lemma 5.3) rather than a direct extension of the $H^{1}$ stability argument used in 1D. The 1D proof seamlessly transitioned from $H^{1}$ stability to uniform $\ell_{\infty}$ boundedness of the inverse operator via the one-dimensional Sobolev embedding theorem, which guarantees that $\left\lVert u\right\rVert_{L^{\infty}(I)}\leq C\left\lVert u\right\rVert_{H^{1}(I)}$. This critical embedding fails for dimensions $d\geq 2$. By proving stability directly in the maximum norm, we bypass the failed embedding and establish the required uniform boundedness needed to prove the exponential decay of the discrete Green’s function.
6. Numerical Validation
We now present numerical experiments to validate the theoretical framework. We analyze the conditioning of the square system ($M=N$) to verify the baseline instability, and we visualize the discrete Green’s function using a stabilized overdetermined solver to confirm the structural dichotomy.
6.1. Baseline Instability
To empirically verify the asymptotic analysis presented in Section 3 and the pre-asymptotic stabilization predicted in Section 4, we analyze the stability of the square system ($M=N$). The numerical experiments were conducted using the basis recombination method to enforce boundary conditions. It is important to note that while recombination algebraically enforces boundary values, it operates within the same underlying Fourier extension frame. Therefore, the asymptotic conditioning properties analyzed in Sections 3 and 4 remain governing factors.
While Theorem 3.1 characterizes the instability via the matrix condition number $\kappa(\bm{A}_{P})$, we verify this experimentally by computing the Lebesgue constant $\Lambda_{N}$. As established in Remark 2.4, $\Lambda_{N}$ acts as a lower bound for the algebraic instability; if $\Lambda_{N}$ grows exponentially, the condition number of the linear system must also grow exponentially. This provides a robust, solver-independent metric for the underlying frame instability.
For the Poisson problem, we set the physical domain length to $L=\pi$ and the exact solution to $u(x)=\sin(\pi x/L)e^{x}$. For the convection-diffusion problem, we set $L=1.0$ with a strong convection coefficient $k=10$, using the exact solution $u(x)=\sin(2\pi x/L)$.
The results, summarized in Figure 3, strikingly confirm the theoretical dichotomy:
  • Poisson Operator (Dashed Lines): As predicted by Theorem 3.1 (via the link in Remark 2.4), the Lebesgue constant $\Lambda_{N}$ grows exponentially with $N$ regardless of the extension parameter $\delta$. While increasing $\delta$ reduces the slope of the growth (improving the constant), it does not alter the fundamental exponential instability. Consequently, the $L^{2}$ error (Figure 3b) hits an accuracy floor (around $10^{-5}$ to $10^{-8}$ for $\delta\leq 1$) and diverges for large $N$. In contrast, the Convection-Diffusion operator delays this saturation significantly. For larger extension parameters (e.g., $\delta=4.0$), it maintains spectral convergence down to machine precision ($\approx 10^{-14}$), whereas the Poisson operator stagnates near $10^{-10}$ even for large $\delta$.
  • Convection-Diffusion Operator (Solid Lines): In stark contrast, the convection-diffusion operator with $k=10$ maintains a lower Lebesgue constant throughout the pre-asymptotic regime. This stability allows the method to achieve near-machine precision accuracy (errors $\approx 10^{-14}$) before the asymptotic frame instability eventually dominates. This validates Corollary 4.8, demonstrating that the quasi-sparse inverse structure effectively suppresses the frame’s ill-conditioning for practical discretizations.
Figure 3. Baseline instability. (a) The Lebesgue constant $\Lambda_{N}$ grows exponentially for both operators, confirming the underlying frame instability. (b) However, the Convection-Diffusion system maintains a Lebesgue constant orders of magnitude lower than Poisson in the pre-asymptotic regime.
6.2. Structural Dichotomy: The Green’s Function
To visualize the intrinsic structure of the inverse operators (Theorem 4.6) independently of the frame instability, we compute the numerical Green’s function. We utilize an overdetermined Least-Squares spectral method with explicit boundary constraints to ensure the solver itself remains robust.
Problem Setting: We solve for the response to a Gaussian source centered at $x_{c}=L/2$ on the physical domain $I=[0,\pi]$:
$\mathcal{L}[u](x)=\exp\left(-\frac{(x-x_{c})^{2}}{2\sigma^{2}}\right),\quad x_{c}=\pi/2,\quad\sigma=0.1.$ (6.1)
The computational domain is extended by $\delta=1.0$. To isolate the structural properties, we employ a stable frame with $N=60$ modes and enforce the boundary conditions $u(0)=u(\pi)=0$ via high-weight constraints in the least-squares system. We compare:
  1. (1)
    Poisson: $\mathcal{L}_{P}=-d^{2}/dx^{2}$.
  2. (2)
    Convection-Diffusion: $\mathcal{L}_{CD}=-d^{2}/dx^{2}+k(d/dx)$, with a moderate convection coefficient $k=5.0$.
Figure 4 plots the computed response $u(x)$.
  • The Poisson response (Blue) is a global, symmetric arch. In the log-scale plot (Panel b), the decay away from the peak is slow and algebraic, indicating a numerically dense inverse matrix.
  • The Convection-Diffusion response (Orange) is an asymmetric, localized spike. Panel (b) reveals a distinct linear slope on the logarithmic scale upstream of the source, confirming the exponential decay predicted by our theory. This ”numerical sparsity” effectively filters the global coupling of errors inherent to the ill-conditioned basis.
Figure 4. Visual verification of the Structural Dichotomy (Theorem 4.6). (a) Linear Scale: The Poisson response (Blue) is a global symmetric arch, indicating that the inverse matrix is dense; information at the center affects the entire domain. The Convection-Diffusion response (Orange) is a localized, asymmetric pulse. (b) Log Scale: The upstream side of the Convection-Diffusion response exhibits a straight linear slope, confirming the exponential decay predicted by our theory. This sparsity isolates numerical errors, explaining the method’s robustness in convection-dominated regimes.
6.3. 2D Validation
We extend the validation to two dimensions to confirm that the structural robustness is preserved in tensor-product domains, as discussed in Section 5. The solver utilizes a tensor-product basis formed by the 1D eigenfunctions, $w_{\mathbf{j}}(\mathbf{x})=w_{j_{1}}(x_{1})w_{j_{2}}(x_{2})$, on the domain $\Omega=[0,1]^{2}$.
We compare three test cases:
  1. (1)
    The Poisson equation $\mathcal{L}_{P}=-\Delta$.
  2. (2)
    Constant-coefficient convection-diffusion with $\mathbf{k}=(10,10)^{T}$.
  3. (3)
    Variable-coefficient convection-diffusion with $\mathbf{k}(x,y)=(5\cos(\pi x),5\cos(\pi y))^{T}$.
Figure 5 presents the $L^{2}$ convergence history. The 2D Poisson operator (Figure 5a) inherits the severe instability of the 1D case; the solver fails to reach high accuracy for small $\delta$ and diverges rapidly as $N$ increases, confirming the compounding instability described in Theorem 5.1.
Conversely, both the constant (Figure 5b) and variable (Figure 5c) coefficient convection-diffusion cases exhibit robust spectral convergence. The stabilizing effect of the convection term is even more pronounced in 2D, consistent with the theoretical requirement of Lemma 5.3, which relies on a sufficiently strong convection field to satisfy the discrete maximum principle. The variable coefficient case demonstrates that this stabilization is robust even when the convection field varies spatially, provided the local Péclet number remains sufficiently high in the pre-asymptotic regime.
Figure 5. 2D Convergence. (a) Poisson. (b) Constant Coeff CD. (c) Variable Coeff CD. The method converges spectrally, demonstrating that the structural robustness extends to tensor-product domains.
7. Conclusion
In this work, we have presented a complete and rigorous stability analysis of the extended-domain spectral collocation method that reconciles its theoretical asymptotic instability with its observed pre-asymptotic behavior. By focusing on the properties of the discrete collocation system and its inverse, we have established a new framework for understanding this widely used method.
Our analysis first confirms the correct asymptotic behavior: the method is fundamentally unstable for both the Poisson and convection-diffusion operators, with a Lebesgue constant that grows exponentially with the number of modes. This finding is rooted in the severe ill-conditioning inherent to the use of a Fourier extension frame on a subinterval. The central contribution of this paper, however, is the explanation for the method’s surprising effectiveness in numerical simulations of convection-dominated problems. We have proven that this is a pre-asymptotic phenomenon, driven by a fundamental difference in the structure of the inverse operators.
We have shown that the transport-dominated nature of the convection-diffusion operator induces a numerically quasi-sparse discrete Green’s function (the inverse of its physical-space representation), whose entries decay exponentially away from the diagonal. In contrast, the global nature of the Poisson operator leads to a numerically dense inverse with slow algebraic decay. For any finite $N$, this structural difference results in a better-conditioned system for the convection-diffusion problem, thereby explaining the numerical observations. These findings provide a solid theoretical foundation for the extended-domain method, giving practitioners a clear understanding of both its practical utility in pre-asymptotic regimes and its ultimate theoretical limitations. Future work could extend this structural analysis of the inverse operator to problems in more complex, non-rectangular geometries.
Acknowledgement
I acknowledge the financial support from the National Science and Technology Council of Taiwan (NSTC 111-2221-E-002-053-MY3).
Appendix A Proof of Lemma 4.1
For completeness, we provide a proof of the refined lemma concerning the exponential decay of the inverse of a matrix with exponentially decaying entries. The proof is an adaptation of the argument presented by Jaffard for infinite-dimensional operators on $\ell^{2}(\mathbb{Z})$ to the case of a sequence of finite-dimensional matrices [jaffard1990proprietes, Prop. 2].
Lemma A.1 (Decay of Inverse Entries).
Let $\{\bm{M}_{N}\}_{N\in\mathbb{N}}$ be a sequence of invertible $N\times N$ matrices. Suppose there exist constants $C_{M}>0$ and $\beta>0$, independent of $N$, such that the entries of each $\bm{M}_{N}$ satisfy
$|(\bm{M}_{N})_{ij}|\leq C_{M}e^{-\beta|i-j|}\quad\text{for all }1\leq i,j\leq N.$
Furthermore, suppose the sequence of inverses is uniformly bounded in the induced $\ell_{\infty}$-norm, i.e., there exists a constant $C_{inv}<\infty$, independent of $N$, such that
$\sup_{N\in\mathbb{N}}\left\lVert\bm{M}_{N}^{-1}\right\rVert_{\infty}\leq C_{inv}.$
Then, there exist constants $C>0$ and $\alpha>0$, independent of $N$, such that the entries of the inverse matrices also decay exponentially:
$|(\bm{M}_{N}^{-1})_{ij}|\leq Ce^{-\alpha|i-j|}\quad\text{for all }1\leq i,j\leq N.$
Proof.
The proof proceeds in five steps. We first establish the necessary operator-norm bounds, then reduce the problem to the positive definite case, and finally apply the core argument from Jaffard [jaffard1990proprietes].
  1. (1)
    Framework and Uniform Boundedness on $\ell^{2}$: Jaffard defines the matrix algebra $\mathcal{E}_{\gamma}$ for matrices whose entries decay as $e^{-\gamma^{\prime}d(s,t)}$ for all $\gamma^{\prime}<\gamma$ [jaffard1990proprietes, Def. 1]. The hypothesis on $\{\bm{M}_{N}\}$ means the sequence is uniformly contained in $\mathcal{E}_{\beta}$.
    By Schur’s Lemma, an operator is bounded on $\ell^{2}$ if its $\ell_{1}$ and $\ell_{\infty}$ induced norms are bounded. The uniform exponential decay of the entries of $\bm{M}_{N}$ implies that the sequence is uniformly bounded in both norms:
    $\left\lVert\bm{M}_{N}\right\rVert_{\infty}=\max_{i}\sum_{j=1}^{N}|(\bm{M}_{N})_{ij}|\leq C_{M}\sum_{k=-\infty}^{\infty}e^{-\beta|k|}=C_{M}\left(1+\frac{2e^{-\beta}}{1-e^{-\beta}}\right)<\infty.$
    The same bound holds for $\left\lVert\bm{M}_{N}\right\rVert_{1}$. Consequently, the $\ell_{2}$-norm is also uniformly bounded: $\left\lVert\bm{M}_{N}\right\rVert_{2}\leq\sqrt{\left\lVert\bm{M}_{N}\right\rVert_{1}\left\lVert\bm{M}_{N}\right\rVert_{\infty}}\leq\text{const}$.
    The proof in [jaffard1990proprietes] assumes invertibility on $\ell^{2}$. The hypothesis $\sup_{N}\left\lVert\bm{M}_{N}^{-1}\right\rVert_{\infty}\leq C_{inv}$ implies that the sequence of operators is uniformly invertible on $\ell^{2}$. This is because any operator bounded on $\ell_{\infty}$ is also bounded on $\ell^{2}$, establishing the necessary condition to apply the logic of the proof.
  2. (2)
    Reduction to the Positive Definite Case: Following Jaffard [jaffard1990proprietes, p. 464], we analyze the inverse by considering the symmetric positive definite matrix $\bm{A}_{N}=\bm{M}_{N}\bm{M}_{N}^{*}$. The inverse of $\bm{M}_{N}$ is then recovered via the identity $\bm{M}_{N}^{-1}=\bm{M}_{N}^{*}\bm{A}_{N}^{-1}$. Since $\mathcal{E}_{\beta}$ is an algebra [jaffard1990proprietes, Prop. 1], and $\{\bm{M}_{N}\},\{\bm{M}_{N}^{*}\}$ are uniformly in $\mathcal{E}_{\beta}$, their product $\{\bm{A}_{N}\}$ is uniformly in $\mathcal{E}_{\beta^{\prime}}$ for any $\beta^{\prime}<\beta$. Furthermore, since $\{\bm{M}_{N}\}$ is uniformly invertible on $\ell^{2}$, so is $\{\bm{A}_{N}\}$. Our goal is to prove that the entries of $\bm{A}_{N}^{-1}$ exhibit uniform exponential decay.
  3. (3)
    Neumann Series Argument: Let $\bm{A}$ be a matrix in the sequence $\{\bm{A}_{N}\}$. We can write it as $\bm{A}=\left\lVert\bm{A}\right\rVert_{2}(\bm{I}-\bm{R})$, where $\bm{R}=\bm{I}-\bm{A}/\left\lVert\bm{A}\right\rVert_{2}$. The uniform invertibility of $\{\bm{A}_{N}\}$ implies that there exists a constant $r<1$, independent of $N$, such that $\left\lVert\bm{R}\right\rVert_{2}\leq r$. The inverse is given by the Neumann series [jaffard1990proprietes, p. 464]:
    $\bm{A}^{-1}=\frac{1}{\left\lVert\bm{A}\right\rVert_{2}}\sum_{k=0}^{\infty}\bm{R}^{k}.$
    The entries of $\bm{R}^{k}$ can be bounded in two ways:
    • (a)
      A norm-based bound: $|(\bm{R}^{k})_{ij}|\leq\left\lVert\bm{R}^{k}\right\rVert_{2}\leq\left\lVert\bm{R}\right\rVert_{2}^{k}\leq r^{k}$.
    • (b)
      An entry-wise decay bound: Since $\bm{R}$ is in $\mathcal{E}_{\beta^{\prime}}$, repeated application of the algebra property shows that for any $\alpha<\beta^{\prime}$, there is a constant $K>1$ (uniform in $N$) such that $|(\bm{R}^{k})_{ij}|\leq K^{k}e^{-\alpha|i-j|}$.
  4. (4)
    Combining Bounds and Summing the Series: Following Jaffard [jaffard1990proprietes, p. 470], we combine these bounds. For any integer $k_{0}>0$, we split the sum:
    $|(\bm{A}^{-1})_{ij}|\leq\frac{1}{\left\lVert\bm{A}\right\rVert_{2}}\left(\sum_{k=0}^{k_{0}}|(\bm{R}^{k})_{ij}|+\sum_{k=k_{0}+1}^{\infty}|(\bm{R}^{k})_{ij}|\right).$
    We bound the first sum using (b) and the second using (a):
    $|(\bm{A}^{-1})_{ij}|\leq\text{const}\cdot\left(\sum_{k=0}^{k_{0}}K^{k}e^{-\alpha|i-j|}+\sum_{k=k_{0}+1}^{\infty}r^{k}\right)$
    $\leq\text{const}\cdot\left((k_{0}+1)K^{k_{0}}e^{-\alpha|i-j|}+\frac{r^{k_{0}+1}}{1-r}\right).$
    We choose $k_{0}$ to balance these two terms. Let $k_{0}=\lceil\frac{\alpha}{\ln(K/r)}|i-j|\rceil$. With this choice, both terms decay exponentially with |ij||i-j|. For instance, the second term becomes:
    $r^{k_{0}}\approx r^{\frac{\alpha\ln(K/r)}{\ln(K/r)}|i-j|}=e^{\ln(r)\frac{\alpha}{\ln(K)-\ln(r)}|i-j|}=e^{-\alpha^{\prime}|i-j|},$
    where $\alpha^{\prime}=\frac{\alpha|\ln(r)|}{\ln(K)-\ln(r)}>0$. The first term exhibits similar behavior. Thus, there exist constants $C^{\prime}>0$ and $\alpha^{\prime}>0$, independent of $N$, such that $|(\bm{A}_{N}^{-1})_{ij}|\leq C^{\prime}e^{-\alpha^{\prime}|i-j|}$.
  5. (5)
    Conclusion: We have shown that $\{\bm{A}_{N}^{-1}\}$ is uniformly in $\mathcal{E}_{\alpha^{\prime}}$. Since $\{\bm{M}_{N}^{*}\}$ is uniformly in $\mathcal{E}_{\beta}$, and $\mathcal{E}$ is an algebra, their product $\bm{M}_{N}^{-1}=\bm{M}_{N}^{*}\bm{A}_{N}^{-1}$ must be uniformly in $\mathcal{E}_{\min(\beta,\alpha^{\prime})}$. This establishes the existence of uniform constants $C$ and $\alpha$ for the decay of entries in $\bm{M}_{N}^{-1}$, completing the proof.
Q.E.D.
Appendix B Technical Lemmas for the Proof of the Maximum Principle
Lemma B.1 (Gårding Inequality with Explicit Constants).
For a continuously differentiable convection field $\mathbf{k}(\mathbf{x})$ and any function $u\in H_{0}^{1}(R_{0})$, the bilinear form $a(u,v)=\int_{R_{0}}\left(\nabla u\cdot\nabla\bar{v}+(\mathbf{k}\cdot\nabla u)\bar{v}\right)\,d\mathbf{x}$ satisfies the Gårding inequality:
$\text{Re}[a(u,u)]\geq\alpha_{0}\left\lVert\nabla u\right\rVert_{L^{2}}^{2}-C_{0}\left\lVert u\right\rVert_{L^{2}}^{2},$
with constants $\alpha_{0}=1$ and $C_{0}=\frac{1}{2}\left\lVert\nabla\cdot\mathbf{k}\right\rVert_{L^{\infty}}$.
Proof.
We analyze the real part of the bilinear form $a(u,u)$ term by term.
$\text{Re}[a(u,u)]=\text{Re}\left[\int_{R_{0}}\left(|\nabla u|^{2}+(\mathbf{k}\cdot\nabla u)\bar{u}\right)\,d\mathbf{x}\right].$
The first term is real and non-negative, giving the coercive part directly:
$\text{Re}\left[\int_{R_{0}}|\nabla u|^{2}\,d\mathbf{x}\right]=\left\lVert\nabla u\right\rVert_{L^{2}}^{2}.$
For the second, convective term, we use the product rule and integration by parts. Consider the divergence of the vector field $\mathbf{F}=\mathbf{k}|u|^{2}=\mathbf{k}(u\bar{u})$:
$\nabla\cdot(\mathbf{k}|u|^{2})=(\nabla\cdot\mathbf{k})|u|^{2}+\mathbf{k}\cdot\nabla(|u|^{2}).$
The gradient of $|u|^{2}$ is $\nabla(u\bar{u})=\bar{u}\nabla u+u\nabla\bar{u}$. Substituting this in gives:
$\nabla\cdot(\mathbf{k}|u|^{2})=(\nabla\cdot\mathbf{k})|u|^{2}+\mathbf{k}\cdot(\bar{u}\nabla u+u\nabla\bar{u}).$
The last term can be written as $\bar{u}(\mathbf{k}\cdot\nabla u)+u(\mathbf{k}\cdot\nabla\bar{u})$. Since $u(\mathbf{k}\cdot\nabla\bar{u})=\overline{\bar{u}(\mathbf{k}\cdot\nabla u)}$, this sum is equal to $2\text{Re}[(\mathbf{k}\cdot\nabla u)\bar{u}]$. Rearranging, we have:
$\text{Re}[(\mathbf{k}\cdot\nabla u)\bar{u}]=\frac{1}{2}\left(\nabla\cdot(\mathbf{k}|u|^{2})-(\nabla\cdot\mathbf{k})|u|^{2}\right).$
Now, we integrate this expression over the domain $R_{0}$:
$\text{Re}\left[\int_{R_{0}}(\mathbf{k}\cdot\nabla u)\bar{u}\,d\mathbf{x}\right]=\frac{1}{2}\int_{R_{0}}\nabla\cdot(\mathbf{k}|u|^{2})\,d\mathbf{x}-\frac{1}{2}\int_{R_{0}}(\nabla\cdot\mathbf{k})|u|^{2}\,d\mathbf{x}.$
By the Divergence Theorem, the first integral on the right becomes a boundary integral: $\int_{\partial R_{0}}(\mathbf{k}|u|^{2})\cdot\mathbf{n}\,dS$. Since $u\in H_{0}^{1}(R_{0})$, $u=0$ on the boundary $\partial R_{0}$, so this boundary integral vanishes. This leaves:
$\text{Re}\left[\int_{R_{0}}(\mathbf{k}\cdot\nabla u)\bar{u}\,d\mathbf{x}\right]=-\frac{1}{2}\int_{R_{0}}(\nabla\cdot\mathbf{k})|u|^{2}\,d\mathbf{x}.$
We can lower-bound this term:
$\frac{1}{2}\int_{R_{0}}(\nabla\cdot\mathbf{k})|u|^{2}\,d\mathbf{x}\geq-\frac{1}{2}\sup_{\mathbf{x}\in R_{0}}|\nabla\cdot\mathbf{k}(\mathbf{x})|\int_{R_{0}}|u|^{2}\,d\mathbf{x}=-\frac{1}{2}\left\lVert\nabla\cdot\mathbf{k}\right\rVert_{L^{\infty}}\left\lVert u\right\rVert_{L^{2}}^{2}.$
Combining the results for the diffusive and convective parts, we get:
$\text{Re}[a(u,u)]\geq\left\lVert\nabla u\right\rVert_{L^{2}}^{2}-\frac{1}{2}\left\lVert\nabla\cdot\mathbf{k}\right\rVert_{L^{\infty}}\left\lVert u\right\rVert_{L^{2}}^{2}.$
This is the Gårding inequality with the identified constants $\alpha_{0}=1$ and $C_{0}=\frac{1}{2}\left\lVert\nabla\cdot\mathbf{k}\right\rVert_{L^{\infty}}$. Q.E.D.
Lemma B.2.
Let $v_{N}\in V_{N,d}$ and let $w_{N}=\mathcal{I}_{N}[v_{N}^{-}]$ be the spectral interpolant of its negative part. The total error between the discrete bilinear form and the continuous energy of the negative part can be bounded for any $\epsilon>0$ by a constant $C_{\epsilon}$ that is independent of $v_{N}$ and $N$ (for $N$ sufficiently large):
$\left|\text{Re}[a_{Q}(v_{N},w_{N})-a(v_{N}^{-},v_{N}^{-})]\right|\leq\epsilon\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}+C_{\epsilon}\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}.$
Proof.
Let $E=\text{Re}[a_{Q}(v_{N},w_{N})-a(v_{N}^{-},v_{N}^{-})]$ be the total error, and let $\eta=v_{N}^{-}-w_{N}$ be the interpolation error of $v_{N}^{-}$. We decompose the error as follows:
$E$ $=\underbrace{\text{Re}[a_{Q}(v_{N},w_{N})-a(v_{N},w_{N})]}_{E_{1}:\text{Quadrature Error}}$
$+\underbrace{\text{Re}[a(v_{N},w_{N})-a(v_{N}^{-},v_{N}^{-})]}_{E_{2}:\text{Interpolation Error}}$
We analyze each term separately, using the following standard properties for the spectral interpolant $\mathcal{I}_{N}$ of a function $f\in H^{1}_{0}(R_{0})$:
  • (P1)
    $L^{2}$-norm error: $\left\lVert f-\mathcal{I}_{N}f\right\rVert_{L^{2}}\leq C_{I}N^{-1}\left\lVert f\right\rVert_{H^{1}}$.
  • (P2)
    $H^{1}$-norm stability: $\left\lVert\nabla(\mathcal{I}_{N}f)\right\rVert_{L^{2}}\leq C_{S}\left\lVert\nabla f\right\rVert_{L^{2}}$.
  • (P3)
    $H^{1}$-orthogonality defect for the Laplacian: $|\langle\nabla f,\nabla(f-\mathcal{I}_{N}f)\rangle|\leq C_{O}N^{-1}\left\lVert f\right\rVert_{H^{1}}^{2}$. This is a known, non-trivial result from spectral approximation theory (see e.g. [bernardi1989properties, Sec. 4]).
1. Bounding $E_{1}$ (Quadrature Error). The term $E_{1}$ is a standard aliasing error for an integrand composed of two polynomials in $V_{N}$. For such terms, established bounds in spectral theory allow us to control the error with lower-order norms. This yields a bound of the required form:
$|E_{1}|\leq\frac{\epsilon}{4}\left\lVert\nabla v_{N}\right\rVert_{H^{1}}^{2}+C_{\epsilon,1}\left\lVert v_{N}\right\rVert_{L^{2}}^{2}.$
Since the supports of $v_{N}^{+}$ and $v_{N}^{-}$ are disjoint, $\left\lVert v_{N}\right\rVert_{H^{1}}^{2}=\left\lVert v_{N}^{+}\right\rVert_{H^{1}}^{2}+\left\lVert v_{N}^{-}\right\rVert_{H^{1}}^{2}$. The term involving $v_{N}^{+}$ can be absorbed into a generic constant, while the term for $v_{N}^{-}$ is kept, leading to:
$|E_{1}|\leq\frac{\epsilon}{4}\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}+C_{\epsilon,1}^{\prime}\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}+\text{terms}(v_{N}^{+}).$
(For simplicity, we will group all terms dependent only on $v_{N}^{+}$ and data into a final constant.)
2. Bounding $E_{2}$ (Interpolation Error). We decompose this term further. Note that $a(v_{N}^{+}+v_{N}^{-},v_{N}^{-})=a(v_{N}^{+},v_{N}^{-})+a(v_{N}^{-},v_{N}^{-})$. Since the supports of $v_{N}^{+}$ and $v_{N}^{-}$ are disjoint, $a(v_{N}^{+},v_{N}^{-})=0$. Thus, $a(v_{N},v_{N}^{-})=a(v_{N}^{-},v_{N}^{-})$. This allows us to simplify $E_{2}$:
$E_{2}=\text{Re}[a(v_{N},w_{N})-a(v_{N},v_{N}^{-})]=-\text{Re}[a(v_{N},\eta)]=-\text{Re}[a(v_{N}^{+},\eta)]-\text{Re}[a(v_{N}^{-},\eta)].$
The term $\text{Re}[a(v_{N}^{+},\eta)]$ can be bounded by standard inequalities and ultimately absorbed into the final constant $C_{\epsilon}$. We focus on the critical term, $\text{Re}[a(v_{N}^{-},\eta)]$.
$a(v_{N}^{-},\eta)=\int_{R_{0}}\left(\nabla v_{N}^{-}\cdot\nabla\bar{\eta}+(\mathbf{k}\cdot\nabla v_{N}^{-})\bar{\eta}\right)d\mathbf{x}$
We bound the two parts of the integrand separately.
  • Diffusion Term: Using the orthogonality defect property (P3) with $f=v_{N}^{-}$:
    $|\langle\nabla v_{N}^{-},\nabla\eta\rangle|\leq C_{O}N^{-1}\left\lVert v_{N}^{-}\right\rVert_{H^{1}}^{2}=C_{O}N^{-1}(\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}+\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}).$
  • Convection Term: Using the Cauchy-Schwarz and Young inequalities, followed by property (P1):
    $|\langle\mathbf{k}\cdot\nabla v_{N}^{-},\eta\rangle|$ $\leq\left\lVert\mathbf{k}\right\rVert_{L^{\infty}}\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}\left\lVert\eta\right\rVert_{L^{2}}$
    $\leq\frac{\epsilon}{4}\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}+\frac{\left\lVert\mathbf{k}\right\rVert_{L^{\infty}}^{2}}{\epsilon}\left\lVert\eta\right\rVert_{L^{2}}^{2}$
    $\leq\frac{\epsilon}{4}\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}+\frac{\left\lVert\mathbf{k}\right\rVert_{L^{\infty}}^{2}C_{I}^{2}N^{-2}}{\epsilon}\left\lVert v_{N}^{-}\right\rVert_{H^{1}}^{2}.$
Combining these bounds for $|a(v_{N}^{-},\eta)|$:
$|a(v_{N}^{-},\eta)|\leq\left(\frac{\epsilon}{4}+C_{O}N^{-1}+\frac{C^{\prime}N^{-2}}{\epsilon}\right)\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}+\left(C_{O}N^{-1}+\frac{C^{\prime}N^{-2}}{\epsilon}\right)\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}.$
For $N$ large enough such that $C_{O}N^{-1}<\epsilon/4$, the coefficient of $\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}$ is less than $\epsilon/2$. The coefficient of $\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}$ is bounded by a constant.
3. Final Assembly. Summing the bounds for all error components ($E_{1}$ and the parts of $E_{2}$), we find that for any chosen $\epsilon>0$, we can choose $N$ large enough such that the total contribution to the $\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}$ term is less than $\epsilon$. All other terms can be collected into a single term $C_{\epsilon}\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}$, where $C_{\epsilon}$ depends on $\epsilon$, operator norms, and interpolation constants, but is independent of $v_{N}$ and $N$. This completes the proof. Q.E.D.